### v2

# Finite-time blowup in a five- and higher-dimensional chemotaxis model with indirect signal production

Abstract.
This paper is concerned with a three-component chemotaxis model accounting for indirect signal production, reading as $u_{t}=\nabla\cdot(\nabla u-u\nabla v)$, $v_{t}=\Delta v-v+w$ and $0=\Delta w-w+u$, posed on a ball in $\mathbb{R}^{n}$ with $n\geq 5$, subject to homogeneous Neumann boundary conditions.
It is suggested that the system has a four-dimensional critical mass phenomenon concerning blowup in finite or infinite time according to the work of Fujie and Senba [J. Differential Equations, 263 (2017), 88–148; 266 (2019), 942–976]. In this paper, we prove that for any prescribed mass $m>0$, there exist radially symmetric and positive initial data $(u_{0},v_{0})\in C^{0}(\overline{\Omega})\times C^{2}(\overline{\Omega})$ with $\int_{\Omega}u_{0}=m$ such that the corresponding solutions blow up in finite time.
Key words and phrases:
chemotaxis; indirect signal production; finite-time blowup
2020 Mathematics Subject Classification:
35B44; 35K51; 35Q92; 92C17
Supported in part by National Natural Science Foundation of China (No. 12271092, No. 11671079) and the Jiangsu Provincial Scientific Research Center of Applied Mathematics (No. BK20233002).
1. Introduction
This paper is concerned with finite-time singularity formation of classical solutions to the following parabolic-parabolic-elliptic chemotaxis model accounting for indirect signal production
$\begin{cases}u_{t}=\Delta u-\nabla\cdot(u\nabla v),&x\in\Omega,t>0,\\ v_{t}=\Delta v-v+w,&x\in\Omega,t>0,\\ 0=\Delta w-w+u,&x\in\Omega,t>0,\\ \partial_{\nu}u=\partial_{\nu}v=\partial_{\nu}w=0,&x\in\partial\Omega,t>0,\\ (u(\cdot,0),v(\cdot,0))=(u_{0},v_{0}),&x\in\Omega,\end{cases}$ (1.1)
in a bounded domain $\Omega\subset\mathbb{R}^{n}$ with smooth boundary for some $n\in\mathbb{N}$, where $\partial_{\nu}$ denotes the derivative with respect to the outward normal of $\partial\Omega$, and the initial functions $u_{0}\in C^{0}(\overline{\Omega})$ and $v_{0}\in C^{2}(\overline{\Omega})$ are nonnegative.
Among the chemotaxis models that describe the oriented movement of organisms in response to chemical stimuli, the minimal chemotaxis system [35]
$\begin{cases}u_{t}=\nabla\cdot(\nabla u-u\nabla v),&x\in\Omega,t>0,\\ v_{t}=\Delta v-v+u,&x\in\Omega,t>0,\end{cases}$ (1.2)
has been extensively studied. This system was originally proposed by Keller and Segel  [19] to investigate the aggregation of cells. The density $u$ of cells is influenced by their motion, which consists of random walk and chemotactic movement along the gradient of chemical concentration $v$. As is well-known, the system exhibits a competitive mechanism between diffusion and aggregation induced by chemotaxis where blowup—the extreme form of aggregation—can occur in finite time under certain conditions:
  • In one dimension, diffusion dominates aggregation, and all solutions of (1.2) are globally bounded [37].
  • In a planar domain, blowup may occur in finite or infinite time if the initial mass $m:=\int_{\Omega}u_{0}$ exceeds the critical mass $4\pi$ but not equals to an integer multiple of $4\pi$ [14]; all solutions remain globally bounded if $m<4\pi$ (or $8\pi$ with symmetry assumptions) [34]. For a disk-shaped domain $\Omega$, there exist radially symmetric initial data with mass $m\in(8\pi,\infty)$ such that the solutions blow up in finite time [31].
  • In higher dimensions, diffusion is rather governed by chemotaxis, and for any initial mass $m>0$, there exist initial data with mass $m$ that lead to finite-time blowup under radial settings [47].
We highlight some related work on (1.2) for interested readers, regarding various aspects: chemotactic collapse [33], asymptotic behaviors under assumptions of uniform-in-time boundedness [9], subcritical mass [44] in two dimensions or smallness conditions [3] in higher dimensions, energy criterion for finite-time blowup [29], blowup rate [30] and blowup profile [48].
Recently, chemotaxis models with indirect signal production, often expressed in the following typical form
$\begin{cases}u_{t}=\Delta u-\nabla\cdot(u\nabla v)+f(u,w),&x\in\Omega,t>0,\\ v_{t}=\Delta v-v+w,&x\in\Omega,t>0,\\ w_{t}=\tau\Delta w-w+u,&x\in\Omega,t>0,\end{cases}$ (1.3)
have received significant attention for their ability to characterize more complex and realistic biological phenomena. For instance, in a study of predicting the cluster attack behaviors of mountain pine beetles [40], the beetle population was divided into two distinct phenotypes: one represented by $u$ for beetles performing chemotaxis and the other by $w$ for beetles producing the attractant $v$, with $f$ modeling potential growth, death and phenotype switching [27]. Here, $\tau\in\{0,1\}$ indicates the motility of the signal producers.
When the signal producer is static, i.e., $\tau=0$, the indirect attractive mechanism has an explosion-suppressing (or boundedness-supporting) effect. Hu and Tao [16] showed that quadratic logistic dampening $f(u,w)=\mu u(1-u)$ to any weak extent $\mu>0$ can prevent blowup in a three-dimensional indirect chemotaxis model with growth (1.3), In contrast, Winkler [46] established boundedness in the three- and higher-dimensional parabolic-parabolic chemotaxis system with a strong logistic source in the sense of $\mu>\mu_{\star}$ for some $\mu_{\star}>0$. For the mass-preserving system (1.3) with $f\equiv 0$, a novel two-dimensional critical mass phenomenon has been identified, concerning infinite-time blowup as documented by Tao and Winkler [42] and Laurençot [21]. For chemotaxis models (1.3) with phenotype switching $f=-u+w$, the occurrence of two-dimensional infinite-time blowup depends on the initial total mass $\int_{\Omega}u_{0}+\int_{\Omega}v_{0}$ [23].
For fully parabolic indirect-chemotaxis models ($\tau=1$) without a source term $f(u,w)\equiv 0$, the indirect attractant mechanism can also reduce explosion-supporting potential. Fujie and Senba [10, 11] established global boundedness in physical spaces $n\leq 3$ and identified a four-dimensional critical mass phenomenon for classical solutions of (1.3). Specifically, they found that
  • Under symmetry assumptions, the solutions exists globally and remain bounded for nonnegative initial data $(u_{0},v_{0},w_{0})$ with $\int_{\Omega}u_{0}<64\pi^{2}$. Without radial symmetry, the same conclusion holds for (1.3) under mixed boundary conditions, where $\partial_{\nu}u-u\partial_{\nu}v=v=w=0$ on $\partial\Omega$ for $t>0$.
  • There exist initial data satisfying $\int_{\Omega}u_{0}\in(64\pi^{2},\infty)\setminus 64\pi^{2}\mathbb{N}$ such that the solutions of (1.3) under mixed boundary conditions blow up in finite or infinite time.
Beyond four dimensions, Mao and Li showed that there exist initial data such that the solutions of the system (1.3) blow up in finite or infinite time [28] with symmetry assumptions. We emphasize both critical dimension $4$ and critical mass $64\pi^{2}$ of (1.3) regarding blowup are significantly larger compared to those of the classical chemotaxis model (1.2). This increase is due to the strong parabolic smoothing effect. Chen and Xiang [5] observed a similar positive effect on boundedness in certain indirect chemotaxis-haptotaxis model. Additionally, taxis-driven aggregation, particularly the spontaneous emergence of large densities during the intermediate transition triggered by a large phenotype switching rate, has been explored in [39, 24].
Concerning finite-time singularity formation, Tao and Winkler [43] considered a Jäger-Luckhaus variant [17] of the system (1.3), where the second equation was modified to $\Delta v-\int_{\Omega}w/|\Omega|+w=0$, and the source term was taken as $f=-au+bw$ for some constants $a\geq 0$ and $b\geq 0$. They showed that there exist radially symmetric initial data such that solutions of the variant, posed on a ball in five and higher dimensions, blow up in finite time. This was deduced by robust comparison methods, which are applicable to cooperative systems involving volume-filling [38] and phenotype switching [27]. In this regard, we focus on detecting finite-time blowup for (1.1), a Nagai-type variant [32] of (1.3). Here, we adopt the assumption that the change of $w$ is stationary over time [18]. However, the comparison arguments used in previous works may not be directly applicable to (1.1) due to the loss of certain partial order relations, inter alia.
Main ideas and results
The system (1.1) has an energy dissipation structure identified by Fujie and Senba [10]. The Lyapunov functional,
$\mathcal{F}(u,v):=\int_{\Omega}(u\log u-uv)+\frac{1}{2}\int_{\Omega}|(-\Delta+1)v|^{2},$
satisfies the differential identity
$\frac{\;\mathrm{d}}{\;\mathrm{d}t}\mathcal{F}(u(t),v(t))+\mathcal{D}(u(t),v(t))=0$
along the reasonably regular solution curves of the system (1.1). Here, the dissipation rate reads as
$\mathcal{D}(u,v):=\int_{\Omega}\left(\left|\nabla(\Delta v-v+w)\right|^{2}+\left|\Delta v-v+w\right|^{2}\right)+\int_{\Omega}u|\nabla(\log u-v)|^{2},$
where $w:=(-\Delta_{N}+1)^{-1}u$ is the unique classical solution of the Neumann elliptic problem
$\begin{cases}-\Delta w+w=u,&x\in\Omega,\\ \partial_{\nu}v=0,&x\in\partial\Omega.\end{cases}$ (1.4)
Invoking the above mathematical structure, we are inspired by previous results (see Remark 1.3) to derive a lower bound for $\mathcal{F}$ in terms of a sublinear power of $\mathcal{D}$ in the form of
$-\mathcal{F}\leq C(\mathcal{D}^{\theta}+1)\quad\text{with some }\theta\in(0,1)\text{ and }C>0,$ (1.5)
for a large class of radially symmetric function pairs $(u,v)=(u(r),v(r))$, which, in particular, includes the radial trajectories of the system (1.1). This result leads to a superlinear differential inequality for $\mathcal{F}$, which, in turn, implies occurrence of finite-time blowup for initial data with low energy.
To be precise, we shall suppose the initial data $(u_{0},v_{0})$ consist of nonnegative functions satisfying
$0\leq u_{0}\in C^{0}(\overline{\Omega}),\quad 0\leq v_{0}\in C^{2}(\overline{\Omega})\quad\text{with }\partial_{\nu}v_{0}\equiv 0\text{ on }\partial\Omega$ (1.6)
and additionally, we impose symmetry assumptions that
$(u_{0},v_{0})\text{ is a pair of positive and radially symmetric functions on }\overline{\Omega},$ (1.7)
with $\Omega=B_{R}:=\{x\in\mathbb{R}^{n}\mid|x|<R\}$ for some $R>0$ and $n\in\mathbb{N}$.
Our main result states that low initial energy enforces finite-time blowup.
Theorem 1.1.
Assume $\Omega=B_{R}\subset\mathbb{R}^{n}$ for some $n\geq 5$ and $R>0$. Let $m>0$ and $A>0$ be given. Then there exist positive constants $K(m,A)$ and $T(m,A)$ such that given any $(u_{0},v_{0})$ from the set
[…]
$\text{and }\mathcal{F}(u_{0},v_{0})<-K(m,A)\},$
for the corresponding classical solution with maximal existence time $T_{\max}(u_{0},v_{0})$ given by Proposition 2.1, one has $T_{\max}(u_{0},v_{0})\leq T(m,A)$.
Concerning the existence of initial data with low energy, for any given positive initial datum with suitable regularity, we construct a family of initial data with large negative energy, which can approximate the given initial datum in an appropriate topology.
Theorem 1.2.
Assume $\Omega=B_{R}\subset\mathbb{R}^{n}$ for some $n\geq 5$ and $R>0$. If $u_{0}\in C^{0}(\overline{\Omega})$ and $v_{0}\in C^{2}(\overline{\Omega})$ comply with (1.6) and (1.7), then one can find a family
$\{(u_{\eta},v_{\eta})\in L^{\infty}(\Omega)\times W^{2,\infty}(\Omega)\}_{\eta\in(0,\eta_{\star})}$
of function couples for some $\eta_{\star}\in(0,1)$ with properties: $(u_{\eta},v_{\eta})$ satisfying (1.6), (1.7) and $\int_{\Omega}u_{\eta}=\int_{\Omega}u_{0}$ for all $\eta\in(0,\eta_{\star})$ is such that
$\|u_{\eta}-u_{0}\|_{L^{p}(\Omega)}+\|v_{\eta}-v_{0}\|_{W^{2,2}(\Omega)}\to 0\quad\text{as }\eta\searrow 0,$
holds for all
$p\in\bigg{[}1,\frac{2n}{n+4}\bigg{)},$
but
$\mathcal{F}(u_{\eta},v_{\eta})\to-\infty\quad\text{as }\eta\searrow 0.$
In particular, for any $\varepsilon>0$, there exist $\eta_{\star\star}=\eta_{\star\star}(\varepsilon)\in(0,\eta_{\star})$, $K(m,A+\varepsilon)$ and $T(m,A+\varepsilon)$ with the property in Theorem 1.1 such that for all $\eta\in(0,\eta_{\star\star})$, $(u_{\eta},v_{\eta})\in\mathcal{B}(m,A+\varepsilon)$ defined as in (1.8), where $\|u_{0}\|_{L^{1}(\Omega)}=m$ and $A=\|v_{0}\|_{W^{2,2}(\Omega)}$.
Remark 1.3.
Our methods are built on the original ideas of Winkler [47], and are also inspired by previous works where functional inequalities of the type in (1.5) have been successfully applied to chemotaxis models in various contexts: volume-filling by Cieślak and Stinner [6, 7, 8], by Laurençot and Mizoguchi [22], by Hashira, Ishida and Yokota [13] and by Cao and Fuest [4]; the Cauchy problem by Winkler [49]; competing chemotaxis by Lankeit [20]; two species models by Li and Li [25] and by Ha and Kim[12]. As far as we are aware, this is the first work to adapt the functional inequality (1.5) to the context of indirect chemotaxis.
The primary challenge lies in deriving the functional inequality (1.5) due to explosion-suppressing effect of indirect attractant mechanisms. Mathematically, we shall obtain interior estimates $\|\Delta v\|_{L^{2}(B_{\rho})}$ in terms of the dissipation rate $\mathcal{D}$ for $\rho\in(0,R)$. To achieve this, we first establish weighted uniform estimates $\||x|^{\beta-1}v\|_{W^{1,\infty}(\Omega)}$ with some $\beta>n-2$ in Section 3. In Section 4, we introduce new techniques to handle the more complex second-order estimate.
This paper is organized as follows. In Section 2, we present some preliminaries. Section 3 is dedicated to deriving uniform-in-time pointwise estimates for the chemical concentration. In Section 4, we establish the functional inequality (1.5). Section 5 is devoted to the proof of Theorem 1.1. In Section 6, we construct a family of positive initial data with arbitrarily large negative energy thus completing the proof of Theorem 1.2.
2. Preliminaries. Energy functional and basic estimates
We begin by revisiting the results on the local existence and uniqueness of classical solutions to the system (1.1).
Proposition 2.1.
Let $n\in\mathbb{N}$ and $\Omega\subset\mathbb{R}^{n}$ be a bounded domain with smooth boundary. Assume that $(u_{0},v_{0})$ is as in (1.6). Then there exist $T_{\max}\in(0,\infty]$ and a unique triplet $(u,v,w)$ of nonnegative functions from $C^{0}\left(\bar{\Omega}\times\left[0,T_{\max}\right)\right)\cap C^{2,1}\left(\bar{\Omega}\times\left(0,T_{\max}\right)\right)$ solving (1.1) classically in $\Omega\times\left(0,T_{\max}\right)$. Moreover $u>0$ in $\Omega\times\left(0,T_{\max}\right)$ and
$\text{if }T_{\max}<\infty,\text{ then }\|u(\cdot,t)\|_{L^{\infty}(\Omega)}\to\infty\text{ as }t\nearrow T_{\max}.$ (2.1)
If in addition $\Omega=B_{R}$ for some $R>0$, and $(u_{0},v_{0})$ is a triplet of radially symmetric functions, then $u$, $v$ and $w$ are all radially symmetric.
Proof.
The local existence, uniqueness and extensibility criterion of classical solutions to the system (1.1) can be established by adapting well-known fixed-point arguments in the contexts of chemotaxis models, as outlined in [15, Theorem 3.1], [41, Lemma 3.1] and [10, Section 4]. The nonnegativity follows from the maximum principle, while conservation of radial symmetry is a consequence of the uniqueness of solutions and the adequate form of the equations in (1.3). ∎
A Lyapunov functional associated with the system (1.1) was formulated in [10, Proposition 6.1].
Proposition 2.2.
Let $(u,v,w)$ be a classical solution given as in Proposition 2.1. Then the following identity holds:
$\frac{d}{dt}\mathcal{F}(u(t),v(t))+\mathcal{D}(u(t),v(t))=0\quad\text{ for all}t\in(0,T_{\max})$ (2.2)
where
$\mathcal{F}(u,v)=\int_{\Omega}(u\log u-uv)+\frac{1}{2}\int_{\Omega}|(-\Delta+1)v|^{2}\quad\text{and}$ (2.3)
$\mathcal{D}(u,v)=\int_{\Omega}\left(\left|\nabla v_{t}\right|^{2}+\left|v_{t}\right|^{2}\right)+\int_{\Omega}u|\nabla(\log u-v)|^{2}.$
For the sake of simplicity in notation, we shall collect some time-independent estimates specifically for high dimensions, i.e., $n\geq 4$. These estimates will serve as a starting point for obtaining certain pointwise estimates in Section 3.
Lemma 2.3.
Let $n\geq 4$ and $(u,v,w)$ be a classical solution given as in Proposition 2.1. Then
$\int_{\Omega}u(x,t)=\int_{\Omega}u_{0}=\int_{\Omega}w(x,t)=:m\quad\text{for all }t\in(0,T_{\max})$ (2.4)
and
$\int_{\Omega}v(x,t)\leq\max\left\{\int_{\Omega}v_{0},\int_{\Omega}u_{0}\right\}\quad\text{for all }t\in(0,T_{\max}).$ (2.5)
For each
$q\in\left[1,\frac{n}{n-1}\right)\quad\text{and}\quad p\in\left[1,\frac{n}{n-2}\right),$
there exists $C_{q}>0$ and $C_{p}>0$ such that
$\|w\|_{W^{1,q}(\Omega)}\leq C_{q}m\quad\text{for all }t\in(0,T_{\max})$ (2.6)
and
$\|v_{t}\|_{L^{p}(\Omega)}+\|w\|_{L^{p}(\Omega)}+\|v\|_{W^{2,p}(\Omega)}\leq C_{p}(m+\|v_{0}\|_{W^{2,2}(\Omega)})\quad\text{for all }t\in(0,T_{\max}),$ (2.7)
respectively.
Proof.
The mass identity (2.4) follows directly from integrating the first and last equations in (1.3) over $\Omega$. Integrating the second equation yields
$\int_{\Omega}v(x,t)=e^{-t}\int_{\Omega}v_{0}+(1-e^{-t})\int_{\Omega}u_{0}\leq\max\left\{\int_{\Omega}v_{0},\int_{\Omega}u_{0}\right\}\quad\text{for all }t\in(0,T_{\max}).$
By applying well-known $W^{1,p}$ regularity theory [1] to the elliptic boundary value problem $-\Delta w+w=u$, we obtain that for each $q\in[1,n/(n-1))$, there exists a constant $C_{q}>0$ such that
$\|w\|_{W^{1,q}(\Omega)}\leq C_{q}\|u\|_{L^{1}(\Omega)}$
holds for all $t\in(0,T_{\max})$. This implies (2.6) in view of (2.4). Thus, $w\in L^{p}(\Omega)$ for $p\in[1,n/(n-2))$ by the Sobolev embedding theorem. Moreover, for any $p\in[1,n/(n-2))$ one can select $q\in[1,n/(n-1))$ such that
$-\frac{1}{2}-\frac{n}{2}\left(\frac{1}{q}-\frac{1}{p}\right)>-1.$ (2.8)
[…]
$\|\nabla v_{x_{i}}\|_{L^{p}(\Omega)}$ $\leq C\|v_{0x_{i}}\|_{W^{1,p}(\Omega)}$
$\quad+C\int_{0}^{t}\left(1+(t-s)^{-\frac{1}{2}-\frac{n}{2}\left(\frac{1}{q}-\frac{1}{p}\right)}\right)e^{-(t-s)}\;\mathrm{d}s\sup_{s\in(0,t)}\|w_{x_{i}}(\cdot,s)\|_{L^{q}(\Omega)}$
$\leq C\|v_{0}\|_{W^{2,2}(\Omega)}+CC_{q}m+CC_{q}m\Gamma\left(\frac{1}{2}-\frac{n}{2}\left(\frac{1}{q}-\frac{1}{p}\right)\right),\quad i\in\{1,2,\cdots,n\},$
holds for all $t\in(0,T_{\max})$, $p\in[1,n/(n-2))$, where the Euler’s Gamma function $\Gamma$ is well-defined due to the choices of parameters in (2.8). According to (2.5), this implies that $v\in W^{2,p}(\Omega)$ for all $t\in(0,T_{\max})$, as well as (2.7), since $v_{t}=\Delta v-v+w\in L^{p}(\Omega)$ for all $t\in(0,T_{\max})$. ∎
3. Pointwise estimates for chemical concentration
This section is devoted to the deduction of uniform-in-time pointwise estimates for $v$ and $w$, based on Lemma 2.3 with additional symmetry assumptions. For convenience, we shall switch to the radial notation without any further comment, e.g., writing $v(r,t)$ instead of $v(x,t)$ when appropriate. Throughout the sequel, we assume that the radius $R$ of the ball $\Omega$ is fixed. Consequently, it will not be explicitly stated that the generic constants $C(p),C(m,M,B,\kappa),\ldots$ may vary from line to line, and may additionally depend on $R$.
The main results of this section are presented as follows.
Proposition 3.1.
Let $\Omega=B_{R}\subset\mathbb{R}^{n}$ with some $R>0$ and $n\geq 4$. If $(u,v,w)$ is a classical solution of (1.1) associated with radially symmetric initial datum $(u_{0},v_{0})$ as given in Proposition 2.1, then there exists a constant $C>0$ such that
$\||x|^{n-1}w\|_{W^{1,\infty}(\Omega)}\leq Cm$ (3.1)
for all $t\in(0,T_{\max})$; and then for each $\beta>n-2$, there exists $C(\beta)>0$ such that
$\||x|^{\beta-1}v\|_{W^{1,\infty}(\Omega)}\leq C(\beta)(m+\|v_{0}\|_{W^{2,2}(\Omega)})$ (3.2)
for all $t\in(0,T_{\max})$.
Proof.
The weighted uniform estimate (3.1) follows directly from Lemma 3.2. The estimates in Lemma 3.3 yield (3.2). ∎
The mass identity (2.4) permits us to gain pointwise information about $w$.
Lemma 3.2.
Let $w$ be as defined in Proposition 3.1. Then there exists a constant $C>0$ such that the nonnegative function $w$ satisfies
$w+|x\cdot\nabla w|\leq\frac{Cm}{|x|^{n-2}},\quad|x|\in(0,R],$ (3.3)
for all $t\in(0,T_{\max})$.
Proof.
Integrating the last equation of (1.1), $ur^{n-1}=wr^{n-1}-(w_{r}r^{n-1})_{r}$ over the interval $(0,\rho)\subset(0,R)$ yields
$|w_{r}|=\left|r^{1-n}\left(\int_{0}^{r}w\eta^{n-1}\;\mathrm{d}\eta-\int_{0}^{r}u\eta^{n-1}\;\mathrm{d}\eta\right)\right|\leq\frac{2\omega_{n}^{-1}m}{r^{n-1}}$
for all $t\in(0,T_{\max})$, where the mass identity (2.4) has been used. Here and henceforth, $\omega_{n}$ denotes the surface area of the unit ball in $\mathbb{R}^{n}$. Choose $r_{0}(t)\in(R/2,R)$ such that
$w(r_{0})\leq\frac{1}{|B_{R}\setminus B_{R/2}|}\int_{B_{R}\setminus B_{R/2}}w\quad\text{for all }t\in(0,T_{\max}).$
Using the Newton-Leibniz formula, we obtain
$w=w(r_{0})+\int_{r_{0}}^{r}w_{r}(\eta)\;\mathrm{d}\eta\leq\frac{m}{|B_{R}\setminus B_{R/2}|}+\frac{2\omega_{n}^{-1}m}{n-2}\left(r^{2-n}+r_{0}^{2-n}\right),$
for all $t\in(0,T_{\max})$. Setting a suitably large constant $C(m)>0$, such as
$C(m):=\frac{m}{|B_{R}\setminus B_{R/2}|}+\frac{2^{n}m}{\omega_{n}},$
we complete the proof. ∎
Following the standard approach in [47], we can deduce pointwise estimates for the gradient of $v$ from (2.7).
Lemma 3.3.
Let $v$ be as defined in Proposition 3.1. Then for any fixed $\beta>n-2$,
$\frac{v}{|x|^{2}}+\frac{|\nabla v|}{|x|}\leq\frac{C(\beta)}{|x|^{\beta}}\left(m+\|v_{0}\|_{W^{2,2}(\Omega)}\right),\quad|x|\in(0,R],\quad t\in(0,T_{\max})$
holds for some $C(\beta)>0$.
Proof.
According to (2.7), we know that $v\in W^{2,p}(\Omega)$ for $p\in[1,n/(n-2))$. Hölder inequality entails
$\left|\int_{r_{1}}^{r}v_{rr}(s)+\frac{n-1}{s}v_{r}(s)\;\mathrm{d}s\right|$
$\leq\left(\int_{0}^{R}|v_{rr}(s)+\frac{n-1}{s}v_{r}(s)|^{p}s^{n-1}\;\mathrm{d}s\right)^{1/p}\left|\int_{r_{1}}^{r}s^{-(n-1)/(p-1)}\;\mathrm{d}s\right|^{\frac{p-1}{p}}$
$\leq\omega_{n}^{-1/p}\|\Delta v\|_{L^{p}(\Omega)}\left(r^{-\frac{n-p}{p}}+r_{1}^{-\frac{n-p}{p}}\right)$
and similarly,
$\left|\int_{r_{1}}^{r}\frac{n-1}{s}v_{r}(s)\;\mathrm{d}s\right|$ $\leq(n-1)\left(\int_{0}^{R}|v_{r}|^{\frac{np}{n-p}}s^{n-1}\;\mathrm{d}s\right)^{\frac{n-p}{np}}\left|\int_{r_{1}}^{r}s^{-\frac{n^{2}-n+p}{np-n+p}}\;\mathrm{d}s\right|^{\frac{np-n+p}{np}}$
$\leq(n-1)\omega_{n}^{-\frac{n-p}{np}}\|\nabla v\|_{L^{\frac{np}{n-p}}(\Omega)}\left(r^{-\frac{n-p}{p}}+r_{1}^{-\frac{n-p}{p}}\right),$
for all $r\in(0,R)$, $p\in(1,n/(n-2))$ and $t\in(0,T_{\max})$, where $r_{1}(t)\in(R/2,R)$ is chosen such that
$|v_{r}(r_{1})|\leq\frac{1}{|B_{R}\setminus B_{R/2}|}\int_{B_{R}\setminus B_{R/2}}|\nabla v|,\quad t\in(0,T_{\max}).$
Newton-Leibniz formula gives
$|v_{r}|$ $=\left|v_{r}(r_{1})+\int_{r_{1}}^{r}v_{rr}(s)\;\mathrm{d}s\right|$
$\leq\left|v_{r}(r_{0})\right|+\left|\int_{r_{1}}^{r}v_{rr}(s)+\frac{n-1}{s}v_{r}(s)\;\mathrm{d}s\right|+\left|\int_{r_{1}}^{r}\frac{n-1}{s}v_{r}(s)\;\mathrm{d}s\right|$
$\leq\frac{\|\nabla v\|_{L^{1}(\Omega)}}{|B_{R}\setminus B_{R/2}|}+C_{s}\|v\|_{W^{2,p}(\Omega)}\left(r^{-\frac{n-p}{p}}+r_{1}^{-\frac{n-p}{p}}\right),$
for all $r\in(0,R)$, $p\in(1,n/(n-2))$ and $t\in(0,T_{\max})$, with some $C_{s}>0$ inferred from the Sobolev inequality. Since $\beta>n-2$, we can put
$p=\max\left\{\frac{n}{n-1},\frac{n}{\beta}\right\}<\frac{n}{n-2}$
and by applying an analogous argument to the one used in Lemma 3.2, we conclude the proof. ∎
4. Energy bounded below by sublinear dissipation rate
The goal of this section is to establish a lower bound for the energy functional $\mathcal{F}$ in terms of a power of the dissipation rate $\mathcal{D}$, as discussed in the introduction, where $\mathcal{F}$ and $\mathcal{D}$ are defined in (2.3).
To relate the energy $\mathcal{F}$ to the dissipation rate $\mathcal{D}$, we introduce the following quantities
$f:=-\Delta v+v-w$ (4.1)
and
$g:=\left(\frac{\nabla u}{\sqrt{u}}-\sqrt{u}\nabla v\right)\cdot\frac{x}{|x|},\quad x\in\overline{\Omega}\setminus\{0\},$ (4.2)
where $w=(-\Delta_{N}+1)^{-1}u$ is the unique classical solution to the Neumann elliptic problem (1.4). Then the energy functional $\mathcal{F}$ and the dissipation rate $\mathcal{D}$ are given by
$\mathcal{F}(u,v)=\int_{\Omega}(u\log u-uv)+\frac{1}{2}\int_{\Omega}|(-\Delta+1)v|^{2}\quad\text{and}$ (4.3)
$\mathcal{D}(u,v)=\int_{\Omega}\left(\left|\nabla f\right|^{2}+\left|f\right|^{2}\right)+\int_{\Omega}g^{2}.$
According to Lemma 2.3 and Proposition 3.1, we will focus on positive and radially symmetric functions $(u,v)$ that satisfy the mass constraints
$\int_{\Omega}u=m,\quad\int_{\Omega}v\leq M\quad\text{and }L^{1}\text{-bound}\quad\int_{\Omega}|f|\leq B,$ (4.4)
as well as the additional pointwise restrictions
$\||x|^{n-1}w\|_{W^{1,\infty}(\Omega)}\leq M\quad\text{and}\quad\||x|^{\kappa-1}v\|_{W^{1,\infty}(\Omega)}\leq B,$ (4.5)
where $m>0$, $M>0$, $B>0$ and $\kappa>n-2$ are fixed parameters. More precisely, the objective is to derive an inequality of the form
$\frac{\mathcal{F}(u,v)}{\mathcal{D}^{\theta}(u,v)+1}\geqslant-C(m,M,B,\kappa)\quad\text{ for all }(u,v)\in\mathcal{S}(m,M,B,\kappa)$
where $\theta\in(0,1)$ and $C(m,M,B,\kappa)>0$. Here $\mathcal{S}(m,M,B,\kappa)$ is the set of functions defined as
[…]
We begin with the following straightforward observation. It can be seen that controlling the energy functional $\mathcal{F}$ from below is equivalent to bounding the mixed term $\int_{\Omega}uv$ from above.
Lemma 4.1.
For all $(u,v)\in\mathcal{S}(m,M,B,\kappa)$,
$-\mathcal{F}(u,v)\leq\int_{\Omega}uv+\frac{\omega_{n}R^{n}}{e}$ (4.6)
holds, where $\omega_{n}$ is the surface area of the unit ball in $\mathbb{R}^{n}$ and $R$ denotes the radius of the ball $\Omega$.
Proof.
This result follows directly from the basic inequality $s\ln s\geq-e^{-1}$ valid for all $s>0$. ∎
Motivated by Winkler [47], we shall estimate the integral $\int_{\Omega}uv$ by separately considering the inner region $B_{\rho}$ and the outer region $\Omega\setminus B_{\rho}$. Additionally, we incorporate decoupling techniques from Cao and Fuest [4], which assist us to convert the mixed term into an interior $L^{2}$–estimate for the Laplacian of the chemical concentration $v$.
Lemma 4.2.
There exists a constant $C(\kappa)>0$ such that
$\int_{\Omega}uv\leq 3\|\Delta v\|_{L^{2}(B_{\rho})}^{2}+3\|v\|_{L^{2}(B_{\rho})}^{2}+\|f\|_{L^{2}(B_{\rho})}^{2}+C(\kappa)(m+M)B\rho^{2-\kappa},$
holds for all $(u,v)\in\mathcal{S}(m,M,B,\kappa)$ and all $\rho\in(0,R)$.
Proof.
For convenience, we introduce a cut-off function defined as
$\phi^{(\rho)}(x):=\phi(x/\rho),\quad x\in\mathbb{R}^{n},\quad\rho\in(0,R),$
where $\phi(x)\in C_{0}^{\infty}(\mathbb{R}^{n})$ is a nonnegative and radially symmetric function such that
$\phi(x)=\begin{cases}1,&\text{if }x\in B_{1/2},\\ 0,&\text{if }x\in\mathbb{R}^{n}\setminus B_{1}.\end{cases}$
It follows that
$\rho\|\nabla\phi^{(\rho)}\|_{L^{\infty}(\mathbb{R}^{n})}+\rho^{2}\|\Delta\phi^{(\rho)}\|_{L^{\infty}(\mathbb{R}^{n})}\leq C_{\phi}:=\|\nabla\phi\|_{L^{\infty}(\mathbb{R}^{n})}+\|\Delta\phi\|_{L^{\infty}(\mathbb{R}^{n})}$
for each $\rho\in(0,R)$. Testing $u=-\Delta w+w$ by $v\phi^{(\rho)}$, we compute upon integration by parts,
$\int_{B_{\rho/2}}uv$ $\leq\int_{\Omega}uv\phi^{(\rho)}=-\int_{\Omega}\Delta wv\phi^{(\rho)}+\int_{\Omega}wv\phi^{(\rho)}=\int_{\Omega}w(-\Delta(v\phi^{(\rho)})+v\phi^{(\rho)})$ (4.7)
$=\int_{\Omega}(-\Delta v+v)w\phi^{(\rho)}-\int_{\Omega}w(2\nabla v\cdot\nabla\phi^{(\rho)}+v\Delta\phi^{(\rho)})=:J_{1}-J_{2}.$
In view of $w=-\Delta v+v-f$ specified in (4.1), we estimate by Young inequality
$J_{1}$ $=\int_{\Omega}(-\Delta v+v)^{2}\phi^{(\rho)}+\int_{\Omega}(\Delta v-v)f\phi^{(\rho)}\leq\frac{3}{2}\|\Delta v-v\|_{L^{2}(B_{\rho})}^{2}+\frac{1}{2}\|f\|_{L^{2}(B_{\rho})}^{2}$ (4.8)
$\leq 3\|\Delta v\|_{L^{2}(B_{\rho})}^{2}+3\|v\|_{L^{2}(B_{\rho})}^{2}+\|f\|_{L^{2}(B_{\rho})}^{2}.$
Recalling pointwise information from (4.5), we estimate
$-J_{2}$ $\leq 2MBC_{\phi}\int_{B_{\rho}\setminus B_{\rho/2}}|x|^{2-n}(|x|^{1-\kappa}|x|^{-1}+|x|^{2-\kappa}|x|^{-2})$ (4.9)
$=2MBC_{\phi}\omega_{n}\int_{\rho/2}^{\rho}r^{1-\kappa}\;\mathrm{d}r\leq 2^{\kappa-1}\omega_{n}MBC_{\phi}\rho^{2-\kappa}.$
Using the pointwise information from (4.5), we also estimate the integral in the annular region
$\int_{\Omega\backslash B_{\rho/2}}uv\leq\sup_{r\in\left(\rho/2,R\right)}v(r)\cdot\int_{\Omega\setminus B_{\rho/2}}u\leq B\sup_{r\in\left(\rho/2,R\right)}r^{2-\kappa}\cdot\int_{\Omega}u=2^{\kappa-2}Bm\rho^{2-\kappa}$ (4.10)
holds for all $(u,v)\in\mathcal{S}$ and $\rho\in(0,R)$. Summing (4.7)–(4.10) yields the desired estimate
$\int_{\Omega}uv\leq 3\|\Delta v\|_{L^{2}(B_{\rho})}^{2}+3\|v\|_{L^{2}(B_{\rho})}^{2}+\|f\|_{L^{2}(B_{\rho})}^{2}+(2^{\kappa-1}\omega_{n}MC_{\phi}+2^{\kappa-2}m)B\rho^{2-\kappa},$
which completes the proof. ∎
To utilize the information from (4.2), we adapt the Pohožaev-type test function $|x|^{n-2}x\cdot\nabla v$, originally proposed by Winkler [45, 46, 47], in order to bound the second-order integral from above by terms from dissipation rate and zeroth-order signal terms. We believe this test scheme has been widely employed in the analysis of elliptic problems [2].
Lemma 4.3.
Let $n\geq 5$. Then there exists a constant $C_{0}>0$ such that for each $\rho\in(0,R)$,
$\frac{1}{8}\|\Delta v\|_{L^{2}(B_{\rho})}^{2}+\frac{3}{4}\|\nabla v\|_{L^{2}(B_{\rho})}^{2}$ $\leq 12\rho^{2}\|\nabla f\|_{L^{2}(B_{\rho})}^{2}+C_{0}\|f\|_{L^{2}(B_{\rho})}^{2}+\sqrt{m}\rho\|g\|_{L^{2}(B_{\rho})}$ (4.11)
$\quad+\|v\|_{L^{2}(B_{\rho})}^{2}+C_{0}B^{2}\rho^{n+2-2\kappa}+C_{0}BM\rho^{2-\kappa}+2m$
holds for all $(u,v)\in\mathcal{S}$.
Proof.
Denote $\alpha=2(n-1)$. Multiplying $\Delta w-w+u=0$ by $|x|^{n-2}x\cdot\nabla v$ and integrating over $B_{\eta}\subset\Omega$ for $\eta\in(0,R)$, we get
$-\int_{0}^{\eta}uv_{r}r^{\alpha}\;\mathrm{d}r+\int_{0}^{\eta}wv_{r}r^{\alpha}\;\mathrm{d}r$ $=\int_{0}^{\eta}(w_{r}r^{n-1})_{r}v_{r}r^{n-1}\;\mathrm{d}r.$ (4.12)
Integration by parts yields
$\int_{0}^{\eta}(w_{r}r^{n-1})_{r}v_{r}r^{n-1}\;\mathrm{d}r$ $=w_{r}(\eta)v_{r}(\eta)\eta^{\alpha}-\int_{0}^{\eta}w_{r}r^{n-1}(v_{r}r^{n-1})_{r}\;\mathrm{d}r$ (4.13)
$=w_{r}(\eta)v_{r}(\eta)\eta^{\alpha}-\int_{0}^{\eta}w_{r}yr^{\alpha}\;\mathrm{d}r,\quad\eta\in(0,R).$
Here, we abbreviate
$y:=\Delta v=v_{rr}+\frac{n-1}{r}v_{r}=(v_{r}r^{n-1})_{r}r^{1-n},\quad r\in(0,R)$
for brevity. Using $w=-y+v-f$ and $w_{r}=-y_{r}+v_{r}-f_{r}$, we evaluate for $\eta\in(0,R)$
$\int_{0}^{\eta}wv_{r}r^{\alpha}\;\mathrm{d}r$ $=\int_{0}^{\eta}(-y+v-f)v_{r}r^{\alpha}\;\mathrm{d}r$ (4.14)
$=-\int_{0}^{\eta}(v_{r}r^{n-1})_{r}v_{r}r^{n-1}\;\mathrm{d}r+\frac{1}{2}\int_{0}^{\eta}(v^{2})_{r}r^{\alpha}\;\mathrm{d}r-\int_{0}^{\eta}fv_{r}r^{\alpha}\;\mathrm{d}r$
$=-\frac{1}{2}v_{r}^{2}(\eta)\eta^{\alpha}+\frac{1}{2}v^{2}(\eta)\eta^{\alpha}-(n-1)\int_{0}^{\eta}v^{2}r^{\alpha-1}\;\mathrm{d}r-\int_{0}^{\eta}fv_{r}r^{\alpha}\;\mathrm{d}r$
and similarly
$-\int_{0}^{\eta}w_{r}yr^{\alpha}\;\mathrm{d}r=-\int_{0}^{\eta}(-y_{r}+v_{r}-f_{r})yr^{\alpha}\;\mathrm{d}r$ (4.15)
$=\int_{0}^{\eta}yy_{r}r^{\alpha}\;\mathrm{d}r-\int_{0}^{\eta}(v_{r}r^{n-1})_{r}v_{r}r^{n-1}\;\mathrm{d}r+\int_{0}^{\eta}f_{r}yr^{\alpha}\;\mathrm{d}r$
$=\frac{1}{2}y^{2}(\eta)\eta^{\alpha}-(n-1)\int_{0}^{\eta}y^{2}r^{\alpha-1}\;\mathrm{d}r-\frac{1}{2}v_{r}^{2}(\eta)\eta^{\alpha}+\int_{0}^{\eta}f_{r}yr^{\alpha}\;\mathrm{d}r.$
Collecting (4.12)–(4.15) and multiplying by $\eta^{1-n}$, we get
$\quad\frac{1}{2}y^{2}(\eta)\eta^{n-1}+w_{r}(\eta)v_{r}(\eta)\eta^{n-1}-\frac{n-1}{\eta^{n-1}}\int_{0}^{\eta}y^{2}r^{\alpha-1}\;\mathrm{d}r$ (4.16)
$=\frac{1}{2}v^{2}(\eta)\eta^{n-1}-\frac{1}{\eta^{n-1}}\int_{0}^{\eta}uv_{r}r^{\alpha}\;\mathrm{d}r-\frac{1}{\eta^{n-1}}\int_{0}^{\eta}fv_{r}r^{\alpha}\;\mathrm{d}r$
$\quad-\frac{1}{\eta^{n-1}}\int_{0}^{\eta}f_{r}yr^{\alpha}\;\mathrm{d}r-\frac{n-1}{\eta^{n-1}}\int_{0}^{\eta}v^{2}r^{\alpha-1}\;\mathrm{d}r$
$\leq\frac{1}{2}v^{2}(\eta)\eta^{n-1}-\frac{1}{\eta^{n-1}}\int_{0}^{\eta}uv_{r}r^{\alpha}\;\mathrm{d}r+\omega_{n}^{-1}\|f\|_{L^{2}(B_{\eta})}\|\nabla v\|_{L^{2}(B_{\eta})}$
$\quad+\omega_{n}^{-1}\|\nabla f\|_{L^{2}(B_{\eta})}\|\Delta v\|_{L^{2}(B_{\eta})}$
for all $\eta\in(0,R)$, where the last inequality follows from dropping the last term and Hölder inequality. Integrating (4.16) over $(0,\rho)\subset(0,R)$, we have
$\frac{1}{2}\int_{0}^{\rho}y^{2}\eta^{n-1}\;\mathrm{d}\eta+\int_{0}^{\rho}w_{r}v_{r}\eta^{n-1}\;\mathrm{d}\eta-\int_{0}^{\rho}\frac{n-1}{\eta^{n-1}}\int_{0}^{\eta}y^{2}r^{\alpha-1}\;\mathrm{d}r\;\mathrm{d}\eta$ (4.17)
$\leq\frac{1}{2}\int_{0}^{\rho}v^{2}\eta^{n-1}\;\mathrm{d}\eta-\int_{0}^{\rho}\frac{1}{\eta^{n-1}}\int_{0}^{\eta}uv_{r}r^{\alpha}\;\mathrm{d}r\;\mathrm{d}\eta$
$\quad+\omega_{n}^{-1}\rho\|\nabla f\|_{L^{2}(B_{\rho})}\|\Delta v\|_{L^{2}(B_{\rho})}+\omega_{n}^{-1}\rho\|f\|_{L^{2}(B_{\rho})}\|\nabla v\|_{L^{2}(B_{\rho})}.$
We calculate the second term on the left of (4.17) by integration by parts, the identity $w=-y+v-f$ from (4.1) and Hölder inequality
$\int_{0}^{\rho}w_{r}v_{r}\eta^{n-1}\;\mathrm{d}\eta=wv_{r}\rho^{n-1}-\int_{0}^{\rho}wy\eta^{n-1}\;\mathrm{d}\eta$ (4.18)
$=wv_{r}\rho^{n-1}+\int_{0}^{\rho}y^{2}\eta^{n-1}\;\mathrm{d}\eta-\int_{0}^{\rho}vy\eta^{n-1}\;\mathrm{d}\eta+\int_{0}^{\rho}fy\eta^{n-1}\;\mathrm{d}\eta$
$=wv_{r}\rho^{n-1}-vv_{r}\rho^{n-1}+\int_{0}^{\rho}y^{2}\eta^{n-1}\;\mathrm{d}\eta+\int_{0}^{\rho}v_{r}^{2}\eta^{n-1}\;\mathrm{d}\eta+\int_{0}^{\rho}fy\eta^{n-1}\;\mathrm{d}\eta$
$\geq\int_{0}^{\rho}y^{2}\eta^{n-1}\;\mathrm{d}\eta+\omega_{n}^{-1}\|\nabla v\|_{L^{2}(B_{\rho})}^{2}-\omega_{n}^{-1}\|\Delta v\|_{L^{2}(B_{\rho})}\|f\|_{L^{2}(B_{\rho})}+(w-v)v_{r}\rho^{n-1}$
for all $\rho\in(0,R)$. By Fubini theorem, we compute the last term on the left of (4.17)
$-\int_{0}^{\rho}\frac{n-1}{\eta^{n-1}}\int_{0}^{\eta}y^{2}r^{\alpha-1}\;\mathrm{d}r\;\mathrm{d}\eta=-(n-1)\int_{0}^{\rho}\int_{r}^{\rho}\frac{\;\mathrm{d}\eta}{\eta^{n-1}}y^{2}r^{\alpha-1}\;\mathrm{d}r$ (4.19)
$=-\frac{n-1}{n-2}\int_{0}^{\rho}y^{2}r^{n-1}\left(1-\frac{r^{n-2}}{\rho^{n-2}}\right)\;\mathrm{d}r\geq-\frac{n-1}{n-2}\int_{0}^{\rho}y^{2}r^{n-1}\;\mathrm{d}r.$
Thank to $uv_{r}=u_{r}-g\sqrt{u}$ as given by (4.2), the second term on the right of (4.17) can be controlled accordingly
$-\int_{0}^{\rho}\frac{1}{\eta^{n-1}}\int_{0}^{\eta}uv_{r}r^{\alpha}\;\mathrm{d}r\;\mathrm{d}\eta=-\int_{0}^{\rho}\frac{1}{\eta^{n-1}}\int_{0}^{\eta}(u_{r}-g\sqrt{u})r^{\alpha}\;\mathrm{d}r\;\mathrm{d}\eta$ (4.20)
$\leq-\int_{0}^{\rho}u\eta^{n-1}\;\mathrm{d}\eta+\alpha\int_{0}^{\rho}\frac{1}{\eta^{n-1}}\int_{0}^{\eta}ur^{\alpha-1}\;\mathrm{d}r\;\mathrm{d}\eta+\int_{0}^{\rho}\int_{0}^{\eta}|g\sqrt{u}|r^{n-1}\;\mathrm{d}r\;\mathrm{d}\eta$
$\leq-\int_{0}^{\rho}u\eta^{n-1}\;\mathrm{d}\eta+\frac{\alpha}{n-2}\int_{0}^{\rho}u\eta^{n-1}\left(1-\frac{\eta^{n-2}}{\rho^{n-2}}\right)\;\mathrm{d}\eta+\frac{\sqrt{m}\rho}{\omega_{n}}\|g\|_{L^{2}(B_{\rho})}$
$\leq\frac{nm}{\omega_{n}(n-2)}+\frac{\sqrt{m}\rho}{\omega_{n}}\|g\|_{L^{2}(B_{\rho})}.$
Substituting (4.18)–(4.20) into (4.17) we get
$\left(\frac{3}{2}-\frac{n-1}{n-2}\right)\|\Delta v\|_{L^{2}(B_{\rho})}^{2}+\|\nabla v\|_{L^{2}(B_{\rho})}^{2}$ (4.21)
$\leq\rho\left(\|\nabla f\|_{L^{2}(B_{\rho})}\|\Delta v\|_{L^{2}(B_{\rho})}+\|\nabla v\|_{L^{2}(B_{\rho})}\|f\|_{L^{2}(B_{\rho})}\right)+\|\Delta v\|_{L^{2}(B_{\rho})}\|f\|_{L^{2}(B_{\rho})}$
$\quad 2m+\sqrt{m}\rho\|g\|_{L^{2}(B_{\rho})}+\|v\|_{L^{2}(B_{\rho})}^{2}+\omega_{n}(v-w)v_{r}\rho^{n-1}.$
The pointwise restrictions (4.5) imply
$\omega_{n}(v-w)v_{r}\rho^{n-1}\leq\omega_{n}(B\rho^{2-\kappa}+M\rho^{2-n})B\rho^{1-\kappa}\rho^{n-1}=\omega_{n}B^{2}\rho^{n+2-2\kappa}+\omega_{n}BM\rho^{2-\kappa}.$
Young inequality gives
$\rho\|\nabla f\|_{L^{2}(B_{\rho})}\|\Delta v\|_{L^{2}(B_{\rho})}\leq\frac{1}{48}\|\Delta v\|_{L^{2}(B_{\rho})}^{2}+12\rho^{2}\|\nabla f\|_{L^{2}(B_{\rho})}^{2},$
and
$\rho\|\nabla v\|_{L^{2}(B_{\rho})}\|f\|_{L^{2}(B_{\rho})}\leq\frac{1}{4}\|\nabla v\|_{L^{2}(B_{\rho})}^{2}+\rho^{2}\|f\|_{L^{2}(B_{\rho})}^{2}$
as well as
$\|\Delta v\|_{L^{2}(B_{\rho})}\|f\|_{L^{2}(B_{\rho})}\leq\frac{1}{48}\|\Delta v\|_{L^{2}(B_{\rho})}+12\|f\|_{L^{2}(B_{\rho})}^{2}.$
Noting the fact
$\frac{3}{2}-\frac{1}{24}-\frac{n-1}{n-2}\geq\frac{1}{8},$
thank to $n\geq 5$, we plug five estimates above into (4.21), which becomes (4.11) by putting $C_{0}:=\max\{\omega_{n},12+R^{2}\}$. ∎
We shall eliminate the zeroth-order terms $\|f\|_{L^{2}(\Omega)}^{2}$ and $\|v\|_{L^{2}(\Omega)}^{2}$ left in Lemma 4.2 and Lemma 4.3 without treatment.
Lemma 4.4.
Let $n\geq 5$. Then there exist $C>0$ and $C(\kappa)>0$ such that
$-\frac{1}{24}\mathcal{F}$ $\leq 12\rho^{2}\|\nabla f\|_{L^{2}(\Omega)}^{2}+CB^{4/(n+2)}\|\nabla f\|_{L^{2}(\Omega)}^{2n/(n+2)}+\sqrt{m}\rho\|g\|_{L^{2}(\Omega)}$
$\quad+C(\kappa)(B^{2}+m^{2}+M^{2}+1)\rho^{n-2\kappa}$
holds for all $(u,v)\in\mathcal{S}$ and $\rho\in(0,R)$, where $\mathcal{F}$ and $\mathcal{D}$ are specified as in (4.3).
Proof.
Lemma 4.1 and 4.2 entail there exists $C(\kappa)>0$ such that
$-\mathcal{F}$ $\leq 3\|\Delta v\|_{L^{2}(B_{\rho})}^{2}+3\|v\|_{L^{2}(B_{\rho})}^{2}+\|f\|_{L^{2}(\Omega)}^{2}+C(\kappa)(Bm+BM+1)\rho^{2-\kappa}$
holds for all $\rho\in(0,R)$. Furthermore, Lemma 4.3 warrants
$-\frac{1}{24}\mathcal{F}$ $\leq 12\rho^{2}\|\nabla f\|_{L^{2}(B_{\rho})}^{2}+C_{0}\|f\|_{L^{2}(B_{\rho})}^{2}+\sqrt{m}\rho\|g\|_{L^{2}(B_{\rho})}$ (4.22)
$\quad+\frac{9}{8}\|v\|_{L^{2}(B_{\rho})}^{2}-\frac{3}{4}\|\nabla v\|_{L^{2}(B_{\rho})}^{2}+C(\kappa)(B^{2}+Bm+BM+m+1)\rho^{n-2\kappa}$
for all $\rho\in(0,R)$. Gagliardo-Nirenberg inequality [36] provides a constant $C_{gn}>0$ such that
$\|\psi\|_{L^{2}(\Omega)}^{2}$ $\leq C_{gn}\|\nabla\psi\|_{L^{2}(\Omega)}^{2n/(n+2)}\|\psi\|_{L^{1}(\Omega)}^{4/(n+2)}+C_{gn}\|\psi\|_{L^{1}(\Omega)}^{2}$
holds for all $\psi\in W^{1,2}(\Omega)$. Recalling (4.4) we thus have
$\|f\|_{L^{2}(\Omega)}^{2}\leq C_{gn}B^{4/(n+2)}\|\nabla f\|_{L^{2}(\Omega)}^{2n/(n+2)}+C_{gn}B^{2},$ (4.23)
as well as
$\|v\|_{L^{2}(B_{\rho})}^{2}$ $\leq C_{gn}M^{4/(n+2)}\|\nabla v\|_{L^{2}(\Omega)}^{2n/(n+2)}+C_{gn}M^{2}$ (4.24)
$\leq\frac{1}{2}\|\nabla v\|_{L^{2}(\Omega)}^{2}+2^{n}C_{gn}^{(n+2)/2}M^{2}+C_{gn}M^{2}$
for all $\rho\in(0,R)$ by Young inequality. Invoking (4.5), we estimate
$\|\nabla v\|^{2}_{L^{2}(\Omega)}$ $=\|\nabla v\|_{L^{2}(B_{\rho})}^{2}+\|\nabla v\|^{2}_{L^{2}(\Omega\setminus B_{\rho})}$ (4.25)
$\leq\|\nabla v\|_{L^{2}(B_{\rho})}^{2}+B^{2}\||x|^{1-\kappa}\|_{L^{2}(\Omega\setminus B_{\rho})}^{2}$
$\leq\|\nabla v\|_{L^{2}(B_{\rho})}^{2}+\omega_{n}B^{2}R^{2}\rho^{n-2\kappa}$
for all $\rho\in(0,R)$. Inserting (4.23)–(4.25) into (4.22) yields
$-\frac{1}{24}\mathcal{F}$ $\leq 12\rho^{2}\|\nabla f\|_{L^{2}(\Omega)}^{2}+C_{0}C_{gn}B^{4/(n+2)}\|\nabla f\|_{L^{2}(\Omega)}^{2n/(n+2)}+\sqrt{m}\rho\|g\|_{L^{2}(\Omega)}$
$\quad+C(\kappa)(B^{2}+Bm+M^{2}+m+1)\rho^{n-2\kappa}$
for all $\rho\in(0,R)$. We finish the proof. ∎
Here we shall reproduce the elementary proof which was given by Winkler [47] to bound the energy $\mathcal{F}$ from below by a sublinear power of the dissipation rate $\mathcal{D}$.
Lemma 4.5.
Let $n\geq 5$. Then there exist $\theta\in(0,1)$ and $C(m,M,B,\kappa)>0$ such that
$-\mathcal{F}\leq C(m,M,B,\kappa)(\mathcal{D}^{\theta}+1)$ (4.26)
holds for all $(u,v)\in\mathcal{S}$, where $\mathcal{F}$ and $\mathcal{D}$ are specified as in (4.3).
Proof.
Put
$\rho:=\min\left\{\frac{R}{2},\|\nabla f\|_{L^{2}(\Omega)}^{-\frac{2n}{(n+2)(2\kappa-n)}}\right\},$
and let
$\theta:=\begin{cases}\frac{n}{n+2}>1-\frac{2n}{(n+2)(2\kappa-n)},&\text{if }\kappa\in(n-2,n),\\ 1-\frac{2n}{(n+2)(2\kappa-n)}\geq\frac{n}{n+2},&\text{if }\kappa\geq n.\end{cases}$
We may infer the existence of $C(m,M,B,\kappa)>0$ from Lemma 4.4 such that
$-\mathcal{F}$ $\leq C(m,M,B,\kappa)(\rho^{2}\|\nabla f\|_{L^{2}(\Omega)}^{2}+\|\nabla f\|_{L^{2}(\Omega)}^{2n/(n+2)}+\|g\|_{L^{2}(\Omega)}+\rho^{n-2\kappa})$
holds for all $(u,v)\in\mathcal{S}$. If $\rho<R/2$, then
$-\mathcal{F}$ $\leq C(m,M,B,\kappa)(\rho^{2}\|\nabla f\|_{L^{2}(\Omega)}^{2}+2\|\nabla f\|_{L^{2}(\Omega)}^{2n/(n+2)}+\|g\|_{L^{2}(\Omega)})$ (4.27)
$\leq C(m,M,B,\kappa)\left(3\|\nabla f\|_{L^{2}(\Omega)}^{2\theta}+\|g\|_{L^{2}(\Omega)}^{2\theta}+3\right)$
$\leq C(m,M,B,\kappa)(\mathcal{D}^{\theta}+1),$
by Young inequality and concavity of the mapping $[0,\infty)\ni\eta\mapsto\eta^{\theta}\in[0,\infty)$. On the other hand, if $\rho=R/2$, then we can estimate analogously
$-\mathcal{F}$ $\leq C(m,M,B,\kappa)\left(\rho^{2}\|\nabla f\|_{L^{2}(\Omega)}^{2}+\|\nabla f\|_{L^{2}(\Omega)}^{2n/(n+2)}+\|g\|_{L^{2}(\Omega)}+\frac{2^{2\kappa-n}}{R^{2\kappa-n}}\right)$ (4.28)
$\leq C(m,M,B,\kappa)\left(2\|\nabla f\|_{L^{2}(\Omega)}^{2\theta}+\|g\|_{L^{2}(\Omega)}^{2\theta}+3+\frac{2^{2\kappa-n}}{R^{2\kappa-n}}\right)$
$\leq C(m,M,B,\kappa)(\mathcal{D}^{\theta}+1).$
Hence, (4.26) follows from (4.27) and (4.28). ∎
5. Finite-time blowup enforced by Low initial energy
This section is devoted to the proof of Theorem 1.1. For initial data $(u_{0},v_{0})$ with low energy $\mathcal{F}(u_{0},v_{0})$, we may deduce a superlinear differential inequality for the mapping $t\mapsto-\mathcal{F}(u(\cdot,t),v(\cdot,t))$ from (4.26), which implies that $(u,v)$ cannot exist globally.
Lemma 5.1.
Assume $\Omega=B_{R}\subset\mathbb{R}^{n}$ for some $n\geq 5$ and $R>0$. Let $m>0$, $A>0$ and $\kappa>n-2$ be given. Then there exist positive constants $K(m,A,\kappa)$ and $T(m,A,\kappa)$ such that given any $(u_{0},v_{0})$ from the set
[…]
$\text{and }\mathcal{F}(u_{0},v_{0})<-K(m,A,\kappa)\},$
for the corresponding classical solution with maximal existence time $T_{\max}(u_{0},v_{0})$ given by Proposition 2.1, one has $T_{\max}(u_{0},v_{0})\leq T(m,A,\kappa)$.
Proof.
Let us fix $c_{1}>0$ such that
$\|\psi\|_{L^{1}(\Omega)}\leq c_{1}\|\psi\|_{W^{2,2}(\Omega)}\quad\text{for all}\psi\in W^{2,2}(\Omega).$
Lemma 2.3 and Proposition 3.1 give $c_{2}>0$, $c_{3}>0$ and $c_{4}=c_{4}(\kappa)>0$ such that whenever $(u_{0},v_{0})$ complies with (1.6), the corresponding solution $(u,v,w)$ satisfies $L^{1}$ estimate
$\|v_{t}\|_{L^{1}(\Omega)}\leq c_{2}(m+\|v_{0}\|_{W^{2,2}(\Omega)})\quad\text{for all }t\in(0,T_{\max}(u_{0},v_{0})),$ (5.2)
as well as pointwise estimates
$\||x|^{n-1}w\|_{W^{1,\infty}(\Omega)}\leq c_{3}m\quad\text{for all }t\in(0,T_{\max}(u_{0},v_{0})),$ (5.3)
and
$\||x|^{\kappa-1}v\|_{W^{1,\infty}(\Omega)}\leq c_{4}(m+\|v_{0}\|_{W^{2,2}(\Omega)})$ (5.4)
for all $t\in(0,T_{\max}(u_{0},v_{0}))$. Now writing
$B$ $:=\max\{c_{1}A,m,c_{3}m\}\quad\text{and}\quad M:=\max\{c_{2}(m+A),c_{4}(m+A)\}$
we invoke Lemma 4.5 to obtain $c_{5}=C(m,M,B,\kappa)>0$ and $\theta\in(0,1)$ such that
$-\mathcal{F}(\tilde{u},\tilde{v})\leq c_{5}(\mathcal{D}^{\theta}(\tilde{u},\tilde{v})+1)\quad\text{for all }(\tilde{u},\tilde{v})\in\mathcal{S}(m,M,B,\kappa).$ (5.5)
We will show the desired conclusion holds if we define
$K(m,A,\kappa):=c_{5}$ (5.6)
and
$T(m,A,\kappa):=\frac{\theta c_{5}^{\frac{1}{\theta}}}{(1-\theta)(-\mathcal{F}(u_{0},v_{0})-c_{5})^{\frac{1-\theta}{\theta}}}.$
Indeed, we know from Proposition 2.1 that for any given $(u_{0},v_{0})\in\widetilde{\mathcal{B}}(m,A,\kappa)$, the corresponding solution $(u,v,w)$ is smooth and radially symmetric with $u>0$ and $v>0$ in $\overline{\Omega}\times[0,T_{\max}(u_{0},v_{0}))$. We infer from (5.2)–(5.4) that
$\|v_{t}\|_{L^{1}(\Omega)}\leq c_{2}(m+A)\leq M\quad\text{for all }t\in(0,T_{\max}(u_{0},v_{0})),$
as well as pointwise estimates
$\||x|^{n-1}w\|_{W^{1,\infty}(\Omega)}\leq c_{3}m\leq B\quad\text{for all }t\in(0,T_{\max}(u_{0},v_{0})),$
and
$\||x|^{\kappa-1}v\|_{W^{1,\infty}(\Omega)}\leq c_{4}(m+A)\leq M$
for all $t\in(0,T_{\max}(u_{0},v_{0}))$. Since moreover $\int_{\Omega}u(x,t)=\int_{\Omega}u_{0}=m$ and
$\int_{\Omega}v(x,t)\leq\max\left\{\int_{\Omega}v_{0},\int_{\Omega}u_{0}\right\}\leq\max\{c_{1}A,m\}\leq B$
for all $t\in(0,T_{\max}(u_{0},v_{0}))$ by (2.4) and (2.5), it follows that $(u,v)\in\mathcal{S}(m,M,B,\kappa)$ for all $t\in(0,T_{\max}(u_{0},v_{0}))$ and thus (5.5) may be applied to $(\tilde{u},\tilde{v})=(u(\cdot,t),v(\cdot,t))$ for all $t\in(0,T_{\max}(u_{0},v_{0}))$. From (2.2) and (5.6), we see that
$\mathcal{F}(u(\cdot,t),v(\cdot,t))\leq\mathcal{F}(u_{0},v_{0})<-K(m,A,\kappa)=-c_{5}\quad\text{for all }t\in(0,T_{\max}(u_{0},v_{0})),$
and thus we may invert (5.5) to get
$-\frac{\;\mathrm{d}}{\;\mathrm{d}t}\mathcal{F}=\mathcal{D}\geq\left(\frac{-\mathcal{F}-c_{5}}{c_{5}}\right)^{\frac{1}{\theta}},\quad\text{for all }t\in(0,T_{\max}(u_{0},v_{0})).$ (5.7)
We shall see $T_{\max}(u_{0},v_{0})\leq T(m,A,\kappa)$. Otherwise a direct comparison argument by solving (5.7) gives
$(-\mathcal{F}(u,v)-c_{5})^{\frac{1-\theta}{\theta}}\geq\frac{\theta c_{5}^{\frac{1}{\theta}}}{(1-\theta)\left(T(m,A,\kappa)-t\right)}$
for all $t\in(0,T(m,A,\kappa))$. This warrants
$\mathcal{F}(u,v)\to-\infty\quad\text{as }t\nearrow T(m,A,\kappa),$
which is incompatible with Proposition 2.1. We thus verify $T_{\max}(u_{0},v_{0})\leq T(m,A,\kappa)$ as desired. ∎
Now we are in a position to show Theorem 1.1.
Proof of Theorem 1.1.
We only need to fix an arbitrary $\kappa>n-2$ and apply Lemma 5.1 to find that the conclusion holds if we let $K(m,A):=K(m,A,\kappa)$ and $T(m,A):=T(m,A,\kappa)$ with $K(m,A,\kappa)$ and $T(m,A,\kappa)$ as given by Lemma 5.1. ∎
6. Initial data with large negative energy
This section is devoted to the existence of initial data with low energy.
Lemma 6.1.
Let $\Omega=B_{R}\subset\mathbb{R}^{n}$ with some $R>0$ and $n\geq 5$. Then for any choice of a pair of positive and radially symmetric functions $(u_{0},v_{0})$ satisfying (1.6), one can find a family
$\{(u_{\eta},v_{\eta})\in L^{\infty}(\Omega)\times W^{2,\infty}(\Omega)\}_{\eta\in(0,\eta_{\star})}$
of function couples for some $\eta_{\star}\in(0,1)$ such that $(u_{\eta},v_{\eta})$ satisfies (1.6) and (1.7) for all $\eta\in(0,\eta_{\star})$, that
$\int_{\Omega}u_{\eta}=\int_{\Omega}u_{0}\quad\text{for all }\eta\in(0,\eta_{\star}),$ (6.1)
that
$v_{\eta}\to v_{0}\quad\text{in }W^{2,2}(\Omega)\text{ as }\eta\searrow 0,$ (6.2)
that
$u_{\eta}\to u_{0}\quad\text{in }L^{p}(\Omega)\text{ for all }p\in\left[1,\frac{2n}{n+4}\right)\text{ as }\eta\searrow 0,$ (6.3)
and that
$\mathcal{F}(u_{\eta},v_{\eta})\to-\infty,\quad\text{as }\eta\searrow 0,$ (6.4)
where $\mathcal{F}$ is taken from (2.3).
Proof.
Let $\phi\in C_{0}^{\infty}(\mathbb{R}^{n})$ be a radially symmetric and nonnegative function with $\int_{\mathbb{R}^{n}}\phi=1$, which is nonincreasing with respect to $|x|$ and supported in $B_{1}$. Fixing $\gamma>1$, we define for $\eta\in(0,\eta_{\star})$,
$u_{\eta}(x):=u_{0}(x)+\left(\ln\frac{1}{\eta}\right)^{2\gamma}\eta^{-\frac{n}{2}-2}\phi\left(\frac{x}{\eta}\right)-\left(\ln\frac{1}{\eta}\right)^{2\gamma}\frac{\eta^{\frac{n}{2}-2}}{|\Omega|},\quad x\in\overline{\Omega},$
and
$v_{\eta}(x):=v_{0}(x)+\left(\ln\frac{1}{\eta}\right)^{-\gamma}\eta^{2-\frac{n}{2}}\phi\left(\frac{x}{\eta}\right),\quad x\in\overline{\Omega},$
where $\eta_{\star}\in(0,1)\cap(0,R)$ is fixed such that
$\eta^{\frac{n}{2}-2}\left(\ln\frac{1}{\eta}\right)^{2\gamma}<|\Omega|\iota\quad\text{with }\iota:=\min_{x\in\overline{\Omega}}u_{0},\quad\text{for all }\eta\in(0,\eta_{\star}),$ (6.5)
and
$2\eta^{\frac{n}{2}-2}\left(\ln\frac{1}{\eta}\right)^{2\gamma}<1\quad\text{for all }\eta\in(0,\eta_{\star}).$ (6.6)
A direct computation can verify the mass identity (6.1). Since
$\|v_{\eta}-v_{0}\|_{W^{2,2}(\Omega)}\leq\left(\ln\frac{1}{\eta}\right)^{-\gamma}\|\phi\|_{W^{2,2}(B_{1})}$
are valid for all $\eta\in(0,\eta_{\star})$, $\gamma>0$ ensures (6.2). We estimate by triangle inequality,
$\|u_{\eta}-u_{0}\|_{L^{p}(\Omega)}\leq\left(\ln\frac{1}{\eta}\right)^{2\gamma}\eta^{-2-\frac{n}{2}+\frac{n}{p}}\|\phi\|_{L^{p}(B_{1})}+\left(\ln\frac{1}{\eta}\right)^{2\gamma}\eta^{\frac{n}{2}-2}|\Omega|^{\frac{1}{p}-1}\quad\text{for all }\eta\in(0,\eta_{\star})$
and (6.3) follows due to $-2-n/2+n/p>0$ and $n>4$. To verify (6.4), we evaluate each term from $\mathcal{F}$. Young inequality entails
$\int_{\Omega}|v_{\eta}|^{2}\leq 2\|v_{0}\|_{L^{2}(\Omega)}^{2}+2\left(\ln\frac{1}{\eta}\right)^{-2\gamma}\eta^{4}\|\phi\|_{L^{2}(B_{1})}^{2}\quad\text{for all }\eta\in(0,\eta_{\star})$ (6.7)
and
$\int_{\Omega}|\Delta v_{\eta}|^{2}\leq 2\|\Delta v_{0}\|_{L^{2}(\Omega)}^{2}+2\left(\ln\frac{1}{\eta}\right)^{-2\gamma}\|\Delta\phi\|_{L^{2}(B_{1})}^{2}\quad\text{for all }\eta\in(0,\eta_{\star}).$ (6.8)
Thank to (6.5), $u_{\eta}$ is nonnegative for all $\eta\in(0,\eta_{\star})$, thus we can drop the nonnegative summands and estimate from below
$\int_{\Omega}u_{\eta}v_{\eta}\geq\left(\ln\frac{1}{\eta}\right)^{\gamma}\int_{\Omega}\phi^{2}\left(\frac{x}{\eta}\right)\eta^{-n}\;\mathrm{d}x=\left(\ln\frac{1}{\eta}\right)^{\gamma}\|\phi\|_{L^{2}(B_{1})}^{2}\quad\text{for all }\eta\in(0,\eta_{\star}).$ (6.9)
Dropping the possible negative part and using the monotonicity of the logarithmic function $\xi\mapsto\ln\xi$ for $\xi>0$, we compute with the help of (6.6)
$\int_{\Omega}u_{\eta}\ln u_{\eta}$ $=\int_{\Omega}u_{\eta}\ln\left(1+\frac{(-\ln\eta)^{2\gamma}\eta^{-\frac{n}{2}-2}}{u_{0}(x)}\phi\left(\frac{x}{\eta}\right)-\frac{(-\ln\eta)^{2\gamma}\eta^{\frac{n}{2}-2}}{u_{0}(x)|\Omega|}\right)+\int_{\Omega}u_{\eta}\ln u_{0}$ (6.10)
$\leq\int_{\Omega}u_{\eta}\ln\left(1+\frac{(-\ln\eta)^{2\gamma}\eta^{-\frac{n}{2}-2}}{u_{0}(x)}\phi\left(\frac{x}{\eta}\right)\right)$
$+\int_{\Omega}u_{0}\ln u_{0}+2\eta^{\frac{n}{2}-2}\left(\ln\frac{1}{\eta}\right)^{2\gamma}\|\ln u_{0}\|_{L^{\infty}(\Omega)}$
$\leq\int_{\Omega}u_{0}\ln u_{0}+\|\ln u_{0}\|_{L^{\infty}(\Omega)}+m\ln\left(1+\left(\ln\frac{1}{\eta}\right)^{2\gamma}\frac{\phi(0)}{\eta^{\frac{n}{2}+2}\iota}\right)$
for all $\eta\in(0,\eta_{\star})$. Using (6.6), we calculate
$\ln\left(1+\left(\ln\frac{1}{\eta}\right)^{2\gamma}\frac{\phi(0)}{\eta^{\frac{n}{2}+2}\iota}\right)$ (6.11)
$=2\gamma\ln\ln\frac{1}{\eta}+\frac{n+4}{2}\ln\frac{1}{\eta}+\ln\frac{\phi(0)}{\iota}+\ln\left(1+\left(\ln\frac{1}{\eta}\right)^{-2\gamma}\frac{\eta^{\frac{n}{2}+2}\iota}{\phi(0)}\right)$
$\leq n\ln\frac{1}{\eta}+\ln\frac{\phi(0)}{\iota}+\ln\left(1+\frac{\iota}{\phi(0)}\right)=n\ln\frac{1}{\eta}+\ln\left(1+\frac{\phi(0)}{\iota}\right)$
for all $\eta\in(0,\eta_{\star})$. Collecting (6.7)–(6.11) gives
$\mathcal{F}(u_{\eta},v_{\eta})=\int_{\Omega}(u_{\eta}\log u_{\eta}-u_{\eta}v_{\eta})+\frac{1}{2}\int_{\Omega}|(-\Delta+1)v_{\eta}|^{2}$ (6.12)
$=-\left(\ln\frac{1}{\eta}\right)^{\gamma}\|\phi\|_{L^{2}(B_{1})}^{2}+2\left(\ln\frac{1}{\eta}\right)^{-2\gamma}(\|\Delta\phi\|_{L^{2}(B_{1})}^{2}+\|\phi\|_{L^{2}(B_{1})}^{2})+nm\ln\frac{1}{\eta}$
$+2\|\Delta v_{0}\|_{L^{2}(\Omega)}^{2}+2\|v_{0}\|_{L^{2}(\Omega)}^{2}+\int_{\Omega}u_{0}\ln u_{0}+\|\ln u_{0}\|_{L^{\infty}(\Omega)}+m\ln\left(1+\frac{\phi(0)}{\iota}\right),$
for all $\eta\in(0,\eta_{\star})$. In view of $\gamma>1$, we have $\mathcal{F}(u_{\eta},v_{\eta})\to-\infty$ as $\eta\searrow 0$ and complete the proof. ∎
Now we are in a position to show Theorem 1.2.
Proof of Theorem 1.2.
Let initial datum $(u_{0},v_{0})$ be specified as in Theorem 1.2. Then the family $\{(u_{\eta},v_{\eta})\}_{\eta\in(0,\eta_{\star})}$ constructed in Lemma 6.1 clearly satisfies two limit formulae according to (6.2)–(6.4).
Let us now fix $\varepsilon>0$, $\|u_{0}\|_{L^{1}(\Omega)}=m$ and $A=\|v_{0}\|_{W^{2,2}(\Omega)}$ given as in Theorem 1.2. We may infer the existence of positive constants $K(m,A+\varepsilon)$ and $T(m,A+\varepsilon)$ from Theorem 1.1 such that for any $(\tilde{u}_{0},\tilde{v}_{0})$ from the set $\mathcal{B}(m,A+\varepsilon)$ defined by (1.8), the corresponding classical solution with maximal existence time $T_{\max}(\tilde{u}_{0},\tilde{v}_{0})$ given by Proposition 2.1, blows up in finite time $T_{\max}(\tilde{u}_{0},\tilde{v}_{0})\leq T(m,A+\varepsilon)$. Thank to (6.1), (6.2) and (6.4), we may pick some $\eta_{\star\star}\in(0,\eta_{\star})$ such that $(u_{\eta},v_{\eta})\in\mathcal{B}(m,A+\varepsilon)$ is valid for all $\eta\in(0,\eta_{\star\star})$. ∎
Acknowledgments
Supported in part by National Natural Science Foundation of China (No. 12271092, No. 11671079) and the Jiangsu Provincial Scientific Research Center of Applied Mathematics (No. BK20233002).