### v3

Abstract
This paper is concerned with a three-component chemotaxis model accounting for indirect signal production, reading as $u_{t}=\nabla\cdot(\nabla u-u\nabla v)$, $v_{t}=\Delta v-v+w$ and $0=\Delta w-w+u$, posed in a ball of $\mathbb{R}^{n}$ with $n\geq 5$, subject to homogeneous Neumann boundary conditions. The system is a Nagai-type variant of its fully parabolic version that has a four-dimensional critical mass phenomenon concerning blowup in finite or infinite time according to the seminal works of Fujie and Senba [J. Differential Equations, 263 (2017), 88–148; 266 (2019), 942–976]. We prove that for any prescribed mass $m>0$, there exist radially symmetric and positive initial data $(u_{0},v_{0})\in C^{0}(\overline{\Omega})\times C^{2}(\overline{\Omega})$ with $\int_{\Omega}u_{0}=m$ such that the corresponding solutions blow up in finite time.
keywords:
chemotaxis, indirect signal production, finite-time blowup
pacs:
MSC Classification]Primary 35B44; Secondary 35K51, 35Q92, 92C17
1 Introduction
This paper is concerned with finite-time singularity formation of classical solutions to the following parabolic-parabolic-elliptic chemotaxis model accounting for indirect signal production
$\begin{cases}u_{t}=\Delta u-\nabla\cdot(u\nabla v),&x\in\Omega,t>0,\\ v_{t}=\Delta v-v+w,&x\in\Omega,t>0,\\ 0=\Delta w-w+u,&x\in\Omega,t>0,\\ \partial_{\nu}u=\partial_{\nu}v=\partial_{\nu}w=0,&x\in\partial\Omega,t>0,\\ (u(\cdot,0),v(\cdot,0))=(u_{0},v_{0}),&x\in\Omega,\end{cases}$ (1)
in a bounded domain $\Omega\subset\mathbb{R}^{n}$ with smooth boundary for some $n\in\mathbb{N}$, where $\partial_{\nu}$ denotes the derivative with respect to the outward normal of $\partial\Omega$, and the initial functions $u_{0}\in C^{0}(\overline{\Omega})$ and $v_{0}\in C^{2}(\overline{\Omega})$ are nonnegative.
Among the chemotaxis models that describe the oriented movement of organisms in response to chemical stimuli, the minimal chemotaxis system
$\begin{cases}u_{t}=\nabla\cdot(\nabla u-u\nabla v),&x\in\Omega,t>0,\\ v_{t}=\Delta v-v+u,&x\in\Omega,t>0,\end{cases}$ (2)
has been extensively studied. This system was originally proposed by Keller and Segel [1] to investigate the aggregation of cells. The density $u$ of cells is influenced by their motion, which consists of random walk and chemotactic movement along the gradient of chemical concentration $v$. As is well-known, the system (2) exhibits a competitive mechanism between diffusion and aggregation induced by chemotaxis where blowup—the extreme form of aggregation—can occur in finite time under certain conditions:
  • In one dimension, diffusion dominates aggregation, and all solutions of (2) are globally bounded [2].
  • In a planar domain, blowup may occur in finite or infinite time if the initial mass $m:=\int_{\Omega}u_{0}$ exceeds the critical mass $4\pi$ but not equals to an integer multiple of $4\pi$ [3]; all solutions remain globally bounded if $m<4\pi$ (or $8\pi$ with symmetry assumptions) [4]. For a disk-shaped domain $\Omega$, there exist radially symmetric initial data with mass $m\in(8\pi,\infty)$ such that the solutions blow up in finite time [5].
  • In higher dimensions, diffusion is rather governed by chemotaxis, and for any initial mass $m>0$, there exist initial data with mass $m$ that lead to finite-time blowup under radial settings [6].
Recently, chemotaxis models involving indirect signal production, often expressed in the following typical form
$\begin{cases}u_{t}=\Delta u-\nabla\cdot(u\nabla v)+f(u,w),&x\in\Omega,t>0,\\ v_{t}=\Delta v-v+w,&x\in\Omega,t>0,\\ w_{t}=\tau\Delta w-w+u,&x\in\Omega,t>0,\end{cases}$ (3)
have received significant attention for their ability to characterize more complex and realistic biological phenomena. For instance, in a study of predicting the cluster attack behaviors of mountain pine beetles [7], the beetle population was divided into two distinct phenotypes: one represented by $u$ for beetles performing chemotaxis and the other by $w$ for beetles producing the attractant $v$, with $f$ modeling potential growth, death and phenotype switching [8]. Here, $\tau\in\{0,1\}$ indicates the motility of the signal producers.
When the signal producer is static, i.e., $\tau=0$, the indirect attractive mechanism has an explosion-suppressing (or boundedness-supporting) effect. Hu and Tao [9] showed that quadratic logistic dampening $f(u,w)=\mu u(1-u)$ to any weak extent $\mu>0$ can prevent blowup in a three-dimensional indirect chemotaxis model with growth (3). In contrast, Winkler [10] established boundedness in the three- and higher-dimensional parabolic-parabolic chemotaxis system with a strong logistic source in the sense of $\mu>\mu_{\star}$ for some $\mu_{\star}>0$. For the mass-preserving system (3) with $f\equiv 0$, a novel two-dimensional critical mass phenomenon has been identified, concerning infinite-time blowup as documented by Tao and Winkler [11] and Laurençot [12]. For chemotaxis models (3) with phenotype switching f=u+wf=-u+w, the occurrence of two-dimensional infinite-time blowup depends on the total mass $\int_{\Omega}u_{0}+\int_{\Omega}v_{0}$ [13].
For fully parabolic indirect-chemotaxis models ($\tau=1$) without a source term $f(u,w)\equiv 0$, the indirect attractant mechanism can also reduce explosion-supporting potential. Fujie and Senba [14, 15] established global boundedness in physical spaces $n\leq 3$ and identified a four-dimensional critical mass phenomenon for classical solutions of (3). Specifically, they found that
  • Under symmetry assumptions, the solutions exists globally and remain bounded for nonnegative initial data $(u_{0},v_{0},w_{0})$ with $\int_{\Omega}u_{0}<64\pi^{2}$. Without radial symmetry, the same conclusion holds for (3) under mixed boundary conditions, where $\partial_{\nu}u-u\partial_{\nu}v=v=w=0$ on $\partial\Omega$ for $t>0$.
  • For any $m\in(64\pi^{2},\infty)\setminus 64\pi^{2}\mathbb{N}$, there exist initial data satisfying $\int_{\Omega}u_{0}=m$ such that the solutions of (3) under mixed boundary conditions blow up in finite or infinite time.
Beyond four dimensions, Mao and Li [16] showed that for any $m>0$, there exist initial data with $\int_{\Omega}u_{0}=m$ such that the solutions of the system (3) blow up in finite or infinite time with symmetry assumptions. We emphasize both critical dimension $4$ and critical mass $64\pi^{2}$ of (3) regarding blowup are significantly larger compared to those of the classical chemotaxis model (2). This increase is due to the strong parabolic smoothing effect of indirect attractant production. Chen and Xiang [17] observed a similar positive effect on boundedness in certain indirect chemotaxis-haptotaxis model. Additionally, taxis-driven aggregation, particularly the spontaneous emergence of large densities during the intermediate transition triggered by large phenotype switching rates, has been explored in [18, 19].
Concerning finite-time singularity formation, Tao and Winkler [20] considered a Jäger-Luckhaus variant of the system (3), where the second equation was modified to $\Delta v-\int_{\Omega}w/|\Omega|+w=0$, and the source term was taken as f=au+bwf=-au+bw for some constants $a\geq 0$ and $b\geq 0$. They showed that there exist radially symmetric initial data such that solutions of the variant, posed in a ball of five and higher dimensions, blow up in finite time. This was deduced by robust comparison methods, which are applicable to cooperative systems involving volume-filling and phenotype switching. In this regard, we focus on detecting finite-time blowup for (1), a Nagai-type variant of (3). Here, we adopt the assumption that the change of $w$ is stationary over time [21]. Notably, the system (3) is also explained as the balanced cases between attraction and repulsion in competing chemotaxis that models the aggregation of microglia observed in Alzheimer’s disease  [22, 14, 15, 23].
Main ideas and results
The system (1) has an energy dissipation structure identified by Fujie and Senba [14]. The Lyapunov functional
$\mathcal{F}(u,v):=\int_{\Omega}(u\log u-uv)+\frac{1}{2}\int_{\Omega}|(-\Delta+1)v|^{2},$
satisfies the differential identity
$\frac{\;\mathrm{d}}{\;\mathrm{d}t}\mathcal{F}(u(t),v(t))+\mathcal{D}(u(t),v(t))=0$
along the reasonably regular solution curves of the system (1). Here, the dissipation rate reads as
$\mathcal{D}(u,v):=\int_{\Omega}\left(\left|\nabla(\Delta v-v+w)\right|^{2}+\left|\Delta v-v+w\right|^{2}\right)+\int_{\Omega}u|\nabla(\log u-v)|^{2},$
where $w:=(-\Delta_{N}+1)^{-1}u$ is the unique classical solution of the Neumann elliptic problem
$\begin{cases}-\Delta w+w=u,&x\in\Omega,\\ \partial_{\nu}w=0,&x\in\partial\Omega.\end{cases}$ (4)
Invoking the above mathematical structure, we are inspired by previous works (see Remark Remark below) to derive a lower bound for $\mathcal{F}$ in terms of a sublinear power of $\mathcal{D}$ in the form of
[…]
for a large class of radially symmetric function pairs $(u,v)=(u(r),v(r))$, which, in particular, includes the radial trajectories of the system (1). This result leads to a superlinear differential inequality for -$\mathcal{F}$, which, in turn, implies occurrence of finite-time blowup for initial data with low energy.
To be precise, we shall suppose the initial data $(u_{0},v_{0})$ consist of nonnegative functions satisfying
$0\lneqq u_{0}\in C^{0}(\overline{\Omega}),\quad 0\leq v_{0}\in C^{2}(\overline{\Omega})\quad\text{with }\partial_{\nu}v_{0}\equiv 0\text{ on }\partial\Omega$ (6)
and additionally, we impose symmetry assumptions that
[…]
with $\Omega=B_{R}:=\{x\in\mathbb{R}^{n}\mid|x|<R\}$ for some $R>0$ and $n\in\mathbb{N}$.
Our main result states that low initial energy enforces finite-time blowup.
Theorem 1.1.
Assume $\Omega=B_{R}\subset\mathbb{R}^{n}$ for some $n\geq 5$ and $R>0$. Let $m>0$ and $A>0$ be given. Then there exist positive constants $K(m,A)$ and $T(m,A)$ such that given any $(u_{0},v_{0})$ from the set
[…]
$\text{ such that }\|u_{0}\|_{L^{1}(\Omega)}=m,\|v_{0}\|_{W^{2,2}(\Omega)}<A\text{ and }\mathcal{F}(u_{0},v_{0})<-K(m,A)\},$
for the corresponding classical solution with maximal existence time $T_{\max}(u_{0},v_{0})$ given by Proposition 2.1, one has $T_{\max}(u_{0},v_{0})\leq T(m,A)$.
Concerning the existence of initial data with low energy, for any given positive initial datum with suitable regularity, we construct a sequence of initial data with arbitrarily large negative energy, which can approximate the given initial datum in an appropriate topology.
Theorem 1.2.
Assume $\Omega=B_{R}\subset\mathbb{R}^{n}$ for some $n\geq 5$ and $R>0$. If $u_{0}\in C^{0}(\overline{\Omega})$ and $v_{0}\in C^{2}(\overline{\Omega})$ comply with (6) and (7), then one can find a sequence
$\{(u_{k},v_{k})\in L^{\infty}(\Omega)\times W^{2,\infty}(\Omega)\}_{k\in\mathbb{N}}$
of function couples with properties: $(u_{k},v_{k})$ satisfying (6), (7) and $\int_{\Omega}u_{k}=\int_{\Omega}u_{0}$ for all $k\in\mathbb{N}$ is such that
$\|u_{k}-u_{0}\|_{L^{p}(\Omega)}+\|v_{k}-v_{0}\|_{W^{2,2}(\Omega)}\to 0\quad\text{as }k\to\infty,$
holds for all $p\in\Big[1,\frac{2n}{n+4}\Big)$, but $\mathcal{F}(u_{k},v_{k})\to$-$\infty$ as $k\to\infty$.
Remark.
Our methods are built on the original ideas of Winkler [6], and are also inspired by previous works where functional inequalities of the type in (5) have been successfully applied to chemotaxis models in various contexts: volume-filling by Cieślak and Stinner [24, 25, 26] and by Cao and Fuest [27]; the Cauchy problem by Winkler [28]; competing chemotaxis by Lankeit [29]; two species models by Li and Li [30]. As far as we are aware, this work presents the first systematic implementation of the functional inequality (5) within the context of indirect chemotaxis.
The primary challenge lies in deriving the functional inequality (5) due to explosion-suppressing effect of indirect attractant mechanisms. Mathematically, we shall obtain interior estimates $\|\Delta v\|_{L^{2}(B_{\rho})}$ in terms of the dissipation rate $\mathcal{D}$ for $\rho\in(0,R)$. To achieve this, we first establish weighted uniform estimates $\||x|^{\beta-1}v\|_{W^{1,\infty}(\Omega)}$ with some $\beta>n-2$ in Section 3. In Section 4, we are stimulated by [6, 31] to introduce the Pohožaev-type test function $\mathds{1}_{B_{\rho}}|x|^{n-2}x\cdot\nabla v$ to obtain the more complex second-order estimate. Here and below, $\mathds{1}_{E}$ stands for the indicator function of a Borel set $E\subseteq\mathbb{R}^{n}$.
This paper is organized as follows. In Section 2, we present some preliminaries involving the energy functional and basic estimates. Section 3 is dedicated to deriving uniform-in-time pointwise estimates for the chemical concentration. In Section 4, we establish the functional inequality (5). Section 5 is devoted to the proof of Theorem 1.1. In Section 6, we construct a sequence of initial data with arbitrarily large negative energy thus completing the proof of Theorem 1.2.
2 Preliminaries
We begin by revisiting the results on the local existence and uniqueness of classical solutions to the system (1).
Proposition 2.1.
Let $n\in\mathbb{N}$ and $\Omega\subset\mathbb{R}^{n}$ be a bounded domain with smooth boundary. Assume that $(u_{0},v_{0})$ is as in (6). Then there exist $T_{\max}\in(0,\infty]$ and a unique triplet $(u,v,w)$ of nonnegative functions from $C^{0}\left(\bar{\Omega}\times\left[0,T_{\max}\right)\right)\cap C^{2,1}\left(\bar{\Omega}\times\left(0,T_{\max}\right)\right)$ solving (1) classically in $\Omega\times\left(0,T_{\max}\right)$. Moreover, $u>0$, $v>0$ and $w>0$ in $\Omega\times\left(0,T_{\max}\right)$, and
$\text{if }T_{\max}<\infty,\text{ then }\|u(\cdot,t)\|_{L^{\infty}(\Omega)}\to\infty\text{ as }t\nearrow T_{\max}.$ (9)
If in addition $\Omega=B_{R}$ for some $R>0$, and $(u_{0},v_{0})$ is a pair of radially symmetric functions, then $u$, $v$ and $w$ are all radially symmetric.
Proof.
The local existence, uniqueness and extensibility criterion of classical solutions to the system (1) can be established by adapting well-known fixed-point arguments in the contexts of chemotaxis models, as outlined in [32, Theorem 3.1], [22, Lemma 3.1] and [14, Section 4]. The nonnegativity follows from the maximum principle, while conservation of radial symmetry is a consequence of the uniqueness of solutions and the adequate form of the equations in (3). ∎
A Lyapunov functional associated with the system (1) was formulated in [14, Proposition 6.1].
Proposition 2.2.
Let $(u,v,w)$ be a classical solution given as in Proposition 2.1. Then the following identity holds:
$\frac{\;\mathrm{d}}{\;\mathrm{d}t}\mathcal{F}(u(t),v(t))+\mathcal{D}(u(t),v(t))=0\quad\text{ for all }t\in(0,T_{\max})$ (10)
where
$\begin{split}&\mathcal{F}(u,v)=\int_{\Omega}(u\log u-uv)+\frac{1}{2}\int_{\Omega}|(-\Delta+1)v|^{2}\quad\text{and}\\ &\mathcal{D}(u,v)=\int_{\Omega}\left(\left|\nabla v_{t}\right|^{2}+\left|v_{t}\right|^{2}\right)+\int_{\Omega}u|\nabla(\log u-v)|^{2}.\end{split}$ (11)
For the sake of simplicity in notation, we shall collect some time-independent estimates specifically for high dimensions, i.e., $n\geq 4$. These estimates will serve as a starting point for obtaining certain pointwise estimates in Section 3.
Lemma 2.3.
Let $\Omega=B_{R}$ for some $R>0$ and $n\geq 4$. Then it holds that for any classical solution $(u,v,w)$ given as in Proposition 2.1,
$\int_{\Omega}u(x,t)=\int_{\Omega}u_{0}=\int_{\Omega}w(x,t)=:m\quad\text{for all }t\in(0,T_{\max})$ (12)
and
$\int_{\Omega}v(x,t)\leq\max\left\{\int_{\Omega}v_{0},\int_{\Omega}u_{0}\right\}\quad\text{for all }t\in(0,T_{\max}).$ (13)
Furthermore, for each $p\in\Big[1,\frac{n}{n-2}\Big)$ and $q\in\Big[1,\frac{n}{n-1}\Big)$, there exist constants $C(p)>0$ and $C(q)>0$ such that for any choice of initial data $(u_{0},v_{0})$ with (6) and (7), the solution of (1) satisfies
$\|w\|_{W^{1,q}(\Omega)}\leq C(q)m\quad\text{for all }t\in(0,T_{\max}),$ (14)
and
$\|v_{t}\|_{L^{p}(\Omega)}+\|w\|_{L^{p}(\Omega)}+\|v\|_{W^{2,p}(\Omega)}\leq C(p)(m+\|v_{0}\|_{W^{2,2}(\Omega)})\quad\text{for all }t\in(0,T_{\max}).$ (15)
Proof.
The mass identity (12) follows directly from integrating the first and last equations in (3) over $\Omega$. Integrating the second equation yields
$\int_{\Omega}v(x,t)=e^{-t}\int_{\Omega}v_{0}+(1-e^{-t})\int_{\Omega}u_{0}\leq\max\left\{\int_{\Omega}v_{0},\int_{\Omega}u_{0}\right\}\quad\text{for all }t\in(0,T_{\max}),$
which shows (13). By applying well-known $W^{1,p}$ regularity theory [33] to the elliptic boundary value problem $\Delta w+w=u$, we obtain that for each $q\in\Big[1,\frac{n}{n-1}\Big)$, there exists a constant $C(q)>0$ such that $\|w\|_{W^{1,q}(\Omega)}\leq C(q)\|u\|_{L^{1}(\Omega)}$ holds for all $t\in(0,T_{\max})$. This implies (14) by virtue of (12). Thus, $w\in L^{p}(\Omega)$ for $p\in\Big[1,\frac{n}{n-2}\Big)$ by Sobolev embedding $W^{1,q}\hookrightarrow L^{\frac{nq}{n-q}}$ for $q\in\Big[1,\frac{n}{n-1}\Big)$. Moreover, for any $p\in\Big[1,\frac{n}{n-2}\Big)$ one can select $q\in\Big[1,\frac{n}{n-1}\Big)$ such that
$\frac{1}{2}-\frac{n}{2}\left(\frac{1}{q}-\frac{1}{p}\right)>-1.$ (16)
$L^{p}$$L^{q}$ estimates of the heat semigroup $e^{t\Delta}$ with homogeneous Dirichlet boundary conditions (cf. [34, Lemma 2.1] and [35, Lemma 2.4]) applied to $v_{x_{i}t}=\Delta v_{x_{i}}-v_{x_{i}}+w_{x_{i}}$ subject to $v_{x_{i}}=v_{r}x_{i}/r\equiv 0$ on $\partial\Omega$ for $i\in\{1,2,\ldots,n\}$, provide a constant $C_{d}>0$ such that
$\|\nabla v_{x_{i}}\|_{L^{p}(\Omega)}$ $\leq\|\nabla e^{t(\Delta-1)}v_{0x_{i}}\|_{L^{p}(\Omega)}+\int_{0}^{t}\|\nabla e^{(t-\tau)(\Delta-1)}w_{x_{i}}\|_{L^{p}(\Omega)}\;\mathrm{d}\tau$
$\leq|\Omega|^{1-p/2}\|\nabla e^{t(\Delta-1)}v_{0x_{i}}\|_{L^{2}(\Omega)}$
$\quad+C_{d}\sup_{\tau\in(0,t)}\|w_{x_{i}}(\cdot,\tau)\|_{L^{q}(\Omega)}\cdot\int_{0}^{t}\left(1+(t-\tau)^{-\frac{1}{2}-\frac{n}{2}\left(\frac{1}{q}-\frac{1}{p}\right)}\right)e^{-(t-\tau)}\;\mathrm{d}\tau$
$\leq C_{d}\|v_{0}\|_{W^{2,2}(\Omega)}+C_{d}C(q)m\bigg(1+\int_{0}^{\infty}\tau^{-\frac{1}{2}-\frac{n}{2}\left(\frac{1}{q}-\frac{1}{p}\right)}e^{-\tau}\;\mathrm{d}\tau\bigg),$
holds for all $t\in(0,T_{\max})$, $p\in\Big[1,\frac{n}{n-2}\Big)$ and $i\in\{1,2,\ldots,n\}$, where the last integral is finite due to the choices of parameters in (16). According to (13), this implies that $v\in W^{2,p}(\Omega)$ for all $t\in(0,T_{\max})$, and (15) holds, since $v_{t}=\Delta v-v+w\in L^{p}(\Omega)$ for all $t\in(0,T_{\max})$. ∎
3 Pointwise estimates for chemical concentration
This section is devoted to the deduction of uniform-in-time pointwise estimates for $v$ and $w$, based on Lemma 2.3. Specifically, we show that symmetric functions from Sobolev spaces can be embedded into certain weighted Sobolev spaces, as established in [36]. For convenience, we shall switch to the radial notation without any further comment, e.g., writing $v(r,t)$ instead of $v(x,t)$ when appropriate. Throughout the sequel, we assume that the radius $R$ of the ball $\Omega$ is fixed. Consequently, it will not be explicitly stated that generic constants $C(p),C(m,M,B,\kappa),\ldots$ may vary from line to line, and additionally depend on $R$.
The main results of this section are presented as follows.
Proposition 3.1.
Let $\Omega=B_{R}\subset\mathbb{R}^{n}$ with some $R>0$ and $n\geq 4$. There exists a constant $C_{w}>0$ such that for any radially symmetric initial datum $(u_{0},v_{0})$ with (6) and (7), the corresponding classical solution $(u,v,w)$ of (1) satisfies
$\sup_{t\in(0,T_{\max})}\||x|^{n-1}w\|_{W^{1,\infty}(\Omega)}\leq C_{w}m.$ (17)
For each $\beta>n-2$, there exists $C(\beta)>0$ such that
$\sup_{t\in(0,T_{\max})}\||x|^{\beta-1}v\|_{W^{1,\infty}(\Omega)}\leq C(\beta)(m+\|v_{0}\|_{W^{2,2}(\Omega)})$ (18)
holds whenever radially symmetric initial data $(u_{0},v_{0})$ comply with (6) and (7).
Proof.
The weighted uniform estimate (17) follows directly from Lemma 3.2. The estimates in Lemma 3.3 yield (18). ∎
The mass identity (12) permits us to gain pointwise information about $w$.
Lemma 3.2.
Let $\Omega=B_{R}\subset\mathbb{R}^{n}$ with some $R>0$ and $n\geq 4$. Then there exists a constant $C_{w}>0$ such that
$w+|x\cdot\nabla w|\leq\frac{C_{w}m}{|x|^{n-2}},\quad|x|\in(0,R],\quad t\in(0,T_{\max})$ (19)
for all $w$ given in Proposition 3.1.
Proof.
Integrating the last equation in (1), $ur^{n-1}=wr^{n-1}-(w_{r}r^{n-1})_{r}$ over the interval $(0,r)\subset(0,R)$ yields
$|w_{r}|=\left|r^{1-n}\left(\int_{0}^{r}w\eta^{n-1}\;\mathrm{d}\eta-\int_{0}^{r}u\eta^{n-1}\;\mathrm{d}\eta\right)\right|\leq\frac{2\omega_{n}^{-1}m}{r^{n-1}}$
for all $t\in(0,T_{\max})$, where the mass identity (12) has been used. Here and henceforth, $\omega_{n}$ denotes the surface area of the unit ball in $\mathbb{R}^{n}$. Choose $r_{0}(t)\in(R/2,R)$ such that
$w(r_{0})\leq\frac{1}{|B_{R}\setminus B_{R/2}|}\int_{B_{R}\setminus B_{R/2}}w\quad\text{for all }t\in(0,T_{\max}).$
Using the Newton-Leibniz formula, we obtain
$\begin{split}w=w(r_{0})+\int_{r_{0}}^{r}w_{r}(\eta)\;\mathrm{d}\eta&\leq\frac{m}{|B_{R}\setminus B_{R/2}|}+\frac{2\omega_{n}^{-1}m}{n-2}\left(r^{2-n}+r_{0}^{2-n}\right)\\ &\leq\frac{mR^{n-2}r^{2-n}}{|B_{R}\setminus B_{R/2}|}+\frac{2\omega_{n}^{-1}m}{n-2}\left(r^{2-n}+2^{n-2}r^{2-n}\right)\end{split}$
for all $t\in(0,T_{\max})$. Setting
$C_{w}:=\frac{R^{n-2}}{|B_{R}\setminus B_{R/2}|}+\frac{2^{n}}{\omega_{n}},$
we complete the proof. ∎
Following the standard approach, e.g. in [37, 6], we can deduce pointwise estimates for the gradient of $v$ from (15). See [36, Theorem 1.1] for generalized results concerning Sobolev imbedding in the spaces of radial functions.
Lemma 3.3.
For any $\beta>n-2$, there exists $C(\beta)>0$ such that
$\frac{v}{|x|^{2}}+\frac{|\nabla v|}{|x|}\leq\frac{C(\beta)}{|x|^{\beta}}\left(m+\|v_{0}\|_{W^{2,2}(\Omega)}\right),\quad|x|\in(0,R],\quad t\in(0,T_{\max})$
holds for all $v$ given in Proposition 3.1.
Proof.
Since $\beta>n-2$, we can put
$p:=\max\bigg\{\frac{n}{n-1},\frac{n}{\beta}\bigg\}\in\bigg(1,\frac{n}{n-2}\bigg).$ (20)
Hölder inequality entails
$\left|\int_{r_{1}}^{r}v_{rr}(s)+\frac{n-1}{s}v_{r}(s)\;\mathrm{d}s\right|$
$\leq\left(\int_{0}^{R}|v_{rr}(s)+\frac{n-1}{s}v_{r}(s)|^{p}s^{n-1}\;\mathrm{d}s\right)^{1/p}\left|\int_{r_{1}}^{r}s^{-(n-1)/(p-1)}\;\mathrm{d}s\right|^{\frac{p-1}{p}}$
$\leq\omega_{n}^{-1/p}\|\Delta v\|_{L^{p}(\Omega)}\left(r^{-\frac{n-p}{p}}+r_{1}^{-\frac{n-p}{p}}\right)$
and similarly,
$\left|\int_{r_{1}}^{r}\frac{n-1}{s}v_{r}(s)\;\mathrm{d}s\right|$ $\leq(n-1)\left(\int_{0}^{R}|v_{r}|^{\frac{np}{n-p}}s^{n-1}\;\mathrm{d}s\right)^{\frac{n-p}{np}}\left|\int_{r_{1}}^{r}s^{-\frac{n^{2}-n+p}{np-n+p}}\;\mathrm{d}s\right|^{\frac{np-n+p}{np}}$
$\leq(n-1)\omega_{n}^{-\frac{n-p}{np}}\|\nabla v\|_{L^{\frac{np}{n-p}}(\Omega)}\left(r^{-\frac{n-p}{p}}+r_{1}^{-\frac{n-p}{p}}\right),$
for all $r\in(0,R)$ and $t\in(0,T_{\max})$, where $r_{1}(t)\in(R/2,R)$ is chosen such that
$|v_{r}(r_{1})|\leq\frac{1}{|B_{R}\setminus B_{R/2}|}\int_{B_{R}\setminus B_{R/2}}|\nabla v|,\quad t\in(0,T_{\max}).$
Newton-Leibniz formula gives
$|v_{r}|$ $=\left|v_{r}(r_{1})+\int_{r_{1}}^{r}v_{rr}(s)\;\mathrm{d}s\right|$
$\leq\left|v_{r}(r_{0})\right|+\left|\int_{r_{1}}^{r}v_{rr}(s)+\frac{n-1}{s}v_{r}(s)\;\mathrm{d}s\right|+\left|\int_{r_{1}}^{r}\frac{n-1}{s}v_{r}(s)\;\mathrm{d}s\right|$
$\leq C_{s}\|v\|_{W^{2,p}(\Omega)}\left(r^{-\frac{n-p}{p}}+r_{1}^{-\frac{n-p}{p}}+\frac{1}{|B_{R}\setminus B_{R/2}|}\right),$
for all $r\in(0,R)$ and $t\in(0,T_{\max})$, with some $C_{s}>0$ inferred from Sobolev imbedding $W^{2,p}(\Omega)\hookrightarrow W^{1,\frac{np}{n-p}}(\Omega)\hookrightarrow W^{1,1}(\Omega)$. According to (15) and (20), the desired pointwise estimates follow from an analogous argument used in Lemma 3.2. ∎
4 Energy bounded below by sublinear dissipation rate
The goal of this section is to establish a lower bound for the energy functional $\mathcal{F}$ in terms of a power of the dissipation rate $\mathcal{D}$, built upon the functional framework originally proposed by Winkler [6], as discussed in the introduction, where $\mathcal{F}$ and $\mathcal{D}$ are defined in (11).
To relate the energy $\mathcal{F}$ to the dissipation rate $\mathcal{D}$ under symmetry assumptions, we introduce the following quantities
$f:=-\Delta v+v-w$ (21)
and
$g:=\left(\frac{\nabla u}{\sqrt{u}}-\sqrt{u}\nabla v\right)\cdot\frac{x}{|x|},\quad x\in\overline{\Omega}\setminus\{0\},$ (22)
where
$w:=(-\Delta_{N}+1)^{-1}u$ (23)
is the unique classical solution to the Neumann elliptic problem (4). Then the energy functional $\mathcal{F}$ and the dissipation rate $\mathcal{D}$ are given by
$\begin{split}&\mathcal{F}(u,v)=\int_{\Omega}(u\log u-uv)+\frac{1}{2}\int_{\Omega}|(-\Delta+1)v|^{2}\quad\text{and}\\ &\mathcal{D}(u,v)=\int_{\Omega}\left(\left|\nabla f\right|^{2}+\left|f\right|^{2}\right)+\int_{\Omega}g^{2}.\end{split}$ (24)
According to Lemma 2.3 and Proposition 3.1, we will focus on positive and radially symmetric functions $(u,v)$ that satisfy the mass constraints
$\int_{\Omega}u=m,\quad\int_{\Omega}v\leq M\quad\text{and }L^{1}\text{-bound}\quad\int_{\Omega}|f|\leq B,$ (25)
as well as the additional pointwise restrictions
$\||x|^{n-1}w\|_{W^{1,\infty}(\Omega)}\leq M\quad\text{and}\quad\||x|^{\kappa-1}v\|_{W^{1,\infty}(\Omega)}\leq B,$ (26)
where $m>0$, $M>0$, $B>0$ and $\kappa>n-2$ are fixed parameters. More precisely, the objective is to derive an inequality of the form
$\frac{\mathcal{F}(u,v)}{\mathcal{D}^{\theta}(u,v)+1}\geqslant-C(m,M,B,\kappa)\quad\text{ for all }(u,v)\in\mathcal{S}(m,M,B,\kappa)$
where $\theta\in(0,1)$ and $C(m,M,B,\kappa)>0$. Here and below, $\mathcal{S}(m,M,B,\kappa)$ is the set of function pairs defined as
$\mathcal{S}(m,M,B,\kappa)$ $=\{(u,v)\in C^{1}(\bar{\Omega})\times C^{3}(\bar{\Omega})\mid u\text{ and }v\text{ are positive and radially }$
[…]
We begin with the following straightforward observation: controlling the energy functional $\mathcal{F}$ from below is equivalent to bounding the mixed term $\int_{\Omega}uv$ from above.
Lemma 4.1.
Let $\Omega=B_{R}\subset\mathbb{R}^{n}$ be a ball of radius $R>0$ with some $n\geq 4$. Then
$\mathcal{F}(u,v)\leq\int_{\Omega}uv+\frac{\omega_{n}R^{n}}{e}$ (27)
holds for all $(u,v)\in\mathcal{S}(m,M,B,\kappa)$, where $\omega_{n}$ is the surface area of the unit ball in $\mathbb{R}^{n}$.
Proof.
It follows from the basic inequality $s\ln s\geq-e^{-1}$ valid for all $s>0$. ∎
We incorporate decoupling techniques from Cao and Fuest [27], which assist us to convert the mixed term $\int_{\Omega}uv$ into an interior $L^{2}$–estimate for the Laplacian of the chemical concentration $v$. We will handle quadratic zeroth-order terms later.
Lemma 4.2.
Let $\Omega=B_{R}\subset\mathbb{R}^{n}$ be a ball of radius $R>0$ with some $n\geq 4$. There exists a constant $C_{2}>0$ such that
$\int_{\Omega}uv\leq 3\|\Delta v\|_{L^{2}(B_{\rho})}^{2}+3\|v\|_{L^{2}(B_{\rho})}^{2}+\|f\|_{L^{2}(B_{\rho})}^{2}+C_{2}(m+M)B\rho^{2-\kappa},$
holds for all $(u,v)\in\mathcal{S}(m,M,B,\kappa)$ and all $\rho\in(0,R)$.
Proof.
In view of (23), we test $u=-\Delta w+w$ by $v\mathds{1}_{B_{\rho}}$, and compute upon integration by parts,
$\begin{split}\int_{B_{\rho}}uv&=-\int_{B_{\rho}}\Delta wv+\int_{B_{\rho}}wv=-\int_{\partial B_{\rho}}v\partial_{\nu}w\;\mathrm{d}S+\int_{B_{\rho}}\nabla w\cdot\nabla v+\int_{B_{\rho}}wv\\ &=\int_{B_{\rho}}(-\Delta v+v)w+\int_{\partial B_{\rho}}(w\partial_{\nu}v-v\partial_{\nu}w)\;\mathrm{d}S=:J_{1}+J_{2}.\end{split}$ (28)
Using $w=-\Delta v+v-f$ specified in (21), we estimate by Young inequality
$\begin{split}J_{1}&=\int_{B_{\rho}}(-\Delta v+v)^{2}+\int_{B_{\rho}}(\Delta v-v)f\leq\frac{3}{2}\|\Delta v-v\|_{L^{2}(B_{\rho})}^{2}+\frac{1}{2}\|f\|_{L^{2}(B_{\rho})}^{2}\\ &\leq 3\|\Delta v\|_{L^{2}(B_{\rho})}^{2}+3\|v\|_{L^{2}(B_{\rho})}^{2}+\|f\|_{L^{2}(B_{\rho})}^{2}.\end{split}$ (29)
Recalling pointwise information from (26), we estimate
$J_{2}\leq MB\int_{\partial B_{\rho}}(|x|^{2-n}|x|^{1-\kappa}+|x|^{2-\kappa}|x|^{1-n})\;\mathrm{d}S=2\omega_{n}MB\rho^{2-\kappa},$ (30)
and also estimate the integral in the annular region
$\int_{\Omega\backslash B_{\rho}}uv\leq\sup_{r\in\left(\rho,R\right)}v(r)\cdot\int_{\Omega\setminus B_{\rho}}u\leq B\sup_{r\in\left(\rho,R\right)}r^{2-\kappa}\cdot\int_{\Omega}u=Bm\rho^{2-\kappa}.$ (31)
Summing (28)–(31) yields the desired estimate
$\int_{\Omega}uv\leq 3\|\Delta v\|_{L^{2}(B_{\rho})}^{2}+3\|v\|_{L^{2}(B_{\rho})}^{2}+\|f\|_{L^{2}(B_{\rho})}^{2}+(2\omega_{n}MB+Bm)\rho^{2-\kappa},$
for all $(u,v)\in\mathcal{S}(m,M,B,\kappa)$ and $\rho\in(0,R)$, which completes the proof. ∎
To leverage the structural information (22), we employ the Pohožaev-type test function $|x|^{n-2}x\cdot\nabla v$ over the inner region $B_{\rho}$, initially introduced in the seminal work of Winkler [38, 10, 6]. This strategy enables us to establish an upper bound for the interior estimate $\|\Delta v\|_{L^{2}(B_{\rho})}$ through a combination of dissipation rate quantities and lower-order signal components. Specifically, this procedure generates a small factor $\rho^{2}$ in the critical term $\rho^{2}\|\nabla f\|_{L^{2}(B_{\rho})}^{2}$. Through careful calibration (see Lemma 4.5), we shall demonstrate that $\rho$ can be adaptively selected depending on $\|\nabla f\|_{L^{2}(\Omega)}$ to effectively convert this quadratic term into a sublinear power of the dissipation rate quantity $\|\nabla f\|_{L^{2}(\Omega)}^{2}$. Such test function methodologies have become standard tools in modern elliptic analysis, as evidenced by recent developments in [31].
Lemma 4.3.
Let $\Omega=B_{R}\subset\mathbb{R}^{n}$ be a ball of radius $R>0$ with some $n\geq 5$. Then there exists a constant $C_{0}>0$ such that
$\begin{split}&\frac{1}{8}\|\Delta v\|_{L^{2}(B_{\rho})}^{2}+\frac{3}{4}\|\nabla v\|_{L^{2}(B_{\rho})}^{2}\leq 12\rho^{2}\|\nabla f\|_{L^{2}(B_{\rho})}^{2}+C_{0}\|f\|_{L^{2}(B_{\rho})}^{2}\\ &+\sqrt{m}\rho\|g\|_{L^{2}(B_{\rho})}+\|v\|_{L^{2}(B_{\rho})}^{2}+C_{0}B^{2}\rho^{n+2-2\kappa}+C_{0}BM\rho^{2-\kappa}+2m\end{split}$ (32)
holds for all $(u,v)\in\mathcal{S}(m,M,B,\kappa)$ and $\rho\in(0,R)$.
Proof.
Denote $\alpha=2(n-1)$. Multiplying $\Delta w-w+u=0$ by $|x|^{n-2}x\cdot\nabla v$ and integrating over $B_{\eta}\subset\Omega$ for $\eta\in(0,R)$, we get
$-\int_{0}^{\eta}uv_{r}r^{\alpha}\;\mathrm{d}r+\int_{0}^{\eta}wv_{r}r^{\alpha}\;\mathrm{d}r$ $=\int_{0}^{\eta}(w_{r}r^{n-1})_{r}v_{r}r^{n-1}\;\mathrm{d}r.$ (33)
Integration by parts yields
$\begin{split}\int_{0}^{\eta}(w_{r}r^{n-1})_{r}v_{r}r^{n-1}\;\mathrm{d}r&=w_{r}(\eta)v_{r}(\eta)\eta^{\alpha}-\int_{0}^{\eta}w_{r}r^{n-1}(v_{r}r^{n-1})_{r}\;\mathrm{d}r\\ &=w_{r}(\eta)v_{r}(\eta)\eta^{\alpha}-\int_{0}^{\eta}w_{r}yr^{\alpha}\;\mathrm{d}r,\quad\eta\in(0,R).\end{split}$ (34)
Here, we abbreviate
$y:=\Delta v=v_{rr}+\frac{n-1}{r}v_{r}=(v_{r}r^{n-1})_{r}r^{1-n},\quad r\in(0,R)$
for brevity. Using w=y+vfw=-y+v-f and $w_{r}=-y_{r}+v_{r}-f_{r}$, we evaluate for $\eta\in(0,R)$
$\begin{split}&\int_{0}^{\eta}wv_{r}r^{\alpha}\;\mathrm{d}r=\int_{0}^{\eta}(-y+v-f)v_{r}r^{\alpha}\;\mathrm{d}r\\ &=-\int_{0}^{\eta}(v_{r}r^{n-1})_{r}v_{r}r^{n-1}\;\mathrm{d}r+\frac{1}{2}\int_{0}^{\eta}(v^{2})_{r}r^{\alpha}\;\mathrm{d}r-\int_{0}^{\eta}fv_{r}r^{\alpha}\;\mathrm{d}r\\ &=-\frac{1}{2}v_{r}^{2}(\eta)\eta^{\alpha}+\frac{1}{2}v^{2}(\eta)\eta^{\alpha}-(n-1)\int_{0}^{\eta}v^{2}r^{\alpha-1}\;\mathrm{d}r-\int_{0}^{\eta}fv_{r}r^{\alpha}\;\mathrm{d}r\end{split}$ (35)
and similarly
$\begin{split}&-\int_{0}^{\eta}w_{r}yr^{\alpha}\;\mathrm{d}r=-\int_{0}^{\eta}(-y_{r}+v_{r}-f_{r})yr^{\alpha}\;\mathrm{d}r\\ &=\int_{0}^{\eta}yy_{r}r^{\alpha}\;\mathrm{d}r-\int_{0}^{\eta}(v_{r}r^{n-1})_{r}v_{r}r^{n-1}\;\mathrm{d}r+\int_{0}^{\eta}f_{r}yr^{\alpha}\;\mathrm{d}r\\ &=\frac{1}{2}y^{2}(\eta)\eta^{\alpha}-(n-1)\int_{0}^{\eta}y^{2}r^{\alpha-1}\;\mathrm{d}r-\frac{1}{2}v_{r}^{2}(\eta)\eta^{\alpha}+\int_{0}^{\eta}f_{r}yr^{\alpha}\;\mathrm{d}r.\end{split}$ (36)
Collecting (33)–(36) and multiplying by $\eta^{1-n}$, we get
$\begin{split}&\frac{1}{2}y^{2}(\eta)\eta^{n-1}+w_{r}(\eta)v_{r}(\eta)\eta^{n-1}-\frac{n-1}{\eta^{n-1}}\int_{0}^{\eta}y^{2}r^{\alpha-1}\;\mathrm{d}r\\ =&\frac{1}{2}v^{2}(\eta)\eta^{n-1}-\frac{1}{\eta^{n-1}}\int_{0}^{\eta}uv_{r}r^{\alpha}\;\mathrm{d}r-\frac{1}{\eta^{n-1}}\int_{0}^{\eta}fv_{r}r^{\alpha}\;\mathrm{d}r\\ &\quad-\frac{1}{\eta^{n-1}}\int_{0}^{\eta}f_{r}yr^{\alpha}\;\mathrm{d}r-\frac{n-1}{\eta^{n-1}}\int_{0}^{\eta}v^{2}r^{\alpha-1}\;\mathrm{d}r\\ \leq&\frac{1}{2}v^{2}(\eta)\eta^{n-1}-\frac{1}{\eta^{n-1}}\int_{0}^{\eta}uv_{r}r^{\alpha}\;\mathrm{d}r+\omega_{n}^{-1}\|f\|_{L^{2}(B_{\eta})}\|\nabla v\|_{L^{2}(B_{\eta})}\\ &\quad+\omega_{n}^{-1}\|\nabla f\|_{L^{2}(B_{\eta})}\|\Delta v\|_{L^{2}(B_{\eta})}\end{split}$ (37)
for all $\eta\in(0,R)$, where the last inequality follows from dropping the last term and Hölder inequality. Integrating (37) over $(0,\rho)\subset(0,R)$, we have
$\begin{split}&\frac{1}{2}\int_{0}^{\rho}y^{2}\eta^{n-1}\;\mathrm{d}\eta+\int_{0}^{\rho}w_{r}v_{r}\eta^{n-1}\;\mathrm{d}\eta-\int_{0}^{\rho}\frac{n-1}{\eta^{n-1}}\int_{0}^{\eta}y^{2}r^{\alpha-1}\;\mathrm{d}r\;\mathrm{d}\eta\\ \leq&\frac{1}{2}\int_{0}^{\rho}v^{2}\eta^{n-1}\;\mathrm{d}\eta-\int_{0}^{\rho}\frac{1}{\eta^{n-1}}\int_{0}^{\eta}uv_{r}r^{\alpha}\;\mathrm{d}r\;\mathrm{d}\eta\\ &\quad+\omega_{n}^{-1}\rho\|\nabla f\|_{L^{2}(B_{\rho})}\|\Delta v\|_{L^{2}(B_{\rho})}+\omega_{n}^{-1}\rho\|f\|_{L^{2}(B_{\rho})}\|\nabla v\|_{L^{2}(B_{\rho})}.\end{split}$ (38)
We calculate the second term on the left of (38) by integration by parts, the identity w=y+vfw=-y+v-f from (21) and Hölder inequality
$\begin{split}&\int_{0}^{\rho}w_{r}v_{r}\eta^{n-1}\;\mathrm{d}\eta=wv_{r}\rho^{n-1}-\int_{0}^{\rho}wy\eta^{n-1}\;\mathrm{d}\eta\\ &=wv_{r}\rho^{n-1}+\int_{0}^{\rho}y^{2}\eta^{n-1}\;\mathrm{d}\eta-\int_{0}^{\rho}vy\eta^{n-1}\;\mathrm{d}\eta+\int_{0}^{\rho}fy\eta^{n-1}\;\mathrm{d}\eta\\ &=wv_{r}\rho^{n-1}-vv_{r}\rho^{n-1}+\int_{0}^{\rho}y^{2}\eta^{n-1}\;\mathrm{d}\eta+\int_{0}^{\rho}v_{r}^{2}\eta^{n-1}\;\mathrm{d}\eta+\int_{0}^{\rho}fy\eta^{n-1}\;\mathrm{d}\eta\\ &\geq\int_{0}^{\rho}y^{2}\eta^{n-1}\;\mathrm{d}\eta+\omega_{n}^{-1}\|\nabla v\|_{L^{2}(B_{\rho})}^{2}-\omega_{n}^{-1}\|\Delta v\|_{L^{2}(B_{\rho})}\|f\|_{L^{2}(B_{\rho})}+(w-v)v_{r}\rho^{n-1}\end{split}$ (39)
for all $\rho\in(0,R)$. By Fubini theorem, we compute the last term on the left of (38)
$\begin{split}&-\int_{0}^{\rho}\frac{n-1}{\eta^{n-1}}\int_{0}^{\eta}y^{2}r^{\alpha-1}\;\mathrm{d}r\;\mathrm{d}\eta=-(n-1)\int_{0}^{\rho}\int_{r}^{\rho}\frac{\;\mathrm{d}\eta}{\eta^{n-1}}y^{2}r^{\alpha-1}\;\mathrm{d}r\\ =&-\frac{n-1}{n-2}\int_{0}^{\rho}y^{2}r^{n-1}\left(1-\frac{r^{n-2}}{\rho^{n-2}}\right)\;\mathrm{d}r\geq-\frac{n-1}{n-2}\int_{0}^{\rho}y^{2}r^{n-1}\;\mathrm{d}r.\end{split}$ (40)
Thanks to $uv_{r}=u_{r}-g\sqrt{u}$ given by (22), the second term on the right of (38) can be controlled accordingly
$\begin{split}&-\int_{0}^{\rho}\frac{1}{\eta^{n-1}}\int_{0}^{\eta}uv_{r}r^{\alpha}\;\mathrm{d}r\;\mathrm{d}\eta=-\int_{0}^{\rho}\frac{1}{\eta^{n-1}}\int_{0}^{\eta}(u_{r}-g\sqrt{u})r^{\alpha}\;\mathrm{d}r\;\mathrm{d}\eta\\ \leq&-\int_{0}^{\rho}u\eta^{n-1}\;\mathrm{d}\eta+\alpha\int_{0}^{\rho}\frac{1}{\eta^{n-1}}\int_{0}^{\eta}ur^{\alpha-1}\;\mathrm{d}r\;\mathrm{d}\eta+\int_{0}^{\rho}\int_{0}^{\eta}|g\sqrt{u}|r^{n-1}\;\mathrm{d}r\;\mathrm{d}\eta\\ \leq&-\int_{0}^{\rho}u\eta^{n-1}\;\mathrm{d}\eta+\frac{\alpha}{n-2}\int_{0}^{\rho}u\eta^{n-1}\left(1-\frac{\eta^{n-2}}{\rho^{n-2}}\right)\;\mathrm{d}\eta+\frac{\sqrt{m}\rho}{\omega_{n}}\|g\|_{L^{2}(B_{\rho})}\\ \leq&\ \frac{nm}{\omega_{n}(n-2)}+\frac{\sqrt{m}\rho}{\omega_{n}}\|g\|_{L^{2}(B_{\rho})}.\end{split}$ (41)
Substituting (39)–(41) into (38) we get
$\begin{split}&\left(\frac{3}{2}-\frac{n-1}{n-2}\right)\|\Delta v\|_{L^{2}(B_{\rho})}^{2}+\|\nabla v\|_{L^{2}(B_{\rho})}^{2}\\ &\leq\rho\left(\|\nabla f\|_{L^{2}(B_{\rho})}\|\Delta v\|_{L^{2}(B_{\rho})}+\|\nabla v\|_{L^{2}(B_{\rho})}\|f\|_{L^{2}(B_{\rho})}\right)+\|\Delta v\|_{L^{2}(B_{\rho})}\|f\|_{L^{2}(B_{\rho})}\\ &\quad 2m+\sqrt{m}\rho\|g\|_{L^{2}(B_{\rho})}+\|v\|_{L^{2}(B_{\rho})}^{2}+\omega_{n}(v-w)v_{r}\rho^{n-1}.\end{split}$ (42)
The pointwise restrictions (26) imply
$\omega_{n}(v-w)v_{r}\rho^{n-1}\leq\omega_{n}(B\rho^{2-\kappa}+M\rho^{2-n})B\rho^{1-\kappa}\rho^{n-1}=\omega_{n}B^{2}\rho^{n+2-2\kappa}+\omega_{n}BM\rho^{2-\kappa}.$
Young inequality gives
$\rho\|\nabla f\|_{L^{2}(B_{\rho})}\|\Delta v\|_{L^{2}(B_{\rho})}\leq\frac{1}{48}\|\Delta v\|_{L^{2}(B_{\rho})}^{2}+12\rho^{2}\|\nabla f\|_{L^{2}(B_{\rho})}^{2},$
and
$\rho\|\nabla v\|_{L^{2}(B_{\rho})}\|f\|_{L^{2}(B_{\rho})}\leq\frac{1}{4}\|\nabla v\|_{L^{2}(B_{\rho})}^{2}+\rho^{2}\|f\|_{L^{2}(B_{\rho})}^{2}$
as well as
$\|\Delta v\|_{L^{2}(B_{\rho})}\|f\|_{L^{2}(B_{\rho})}\leq\frac{1}{48}\|\Delta v\|_{L^{2}(B_{\rho})}+12\|f\|_{L^{2}(B_{\rho})}^{2}.$
Noting the fact
$\frac{3}{2}-\frac{1}{24}-\frac{n-1}{n-2}\geq\frac{1}{8},$
thanks to $n\geq 5$, we plug five estimates above into (42), which becomes (32) by putting $C_{0}:=\max\{\omega_{n},12+R^{2}\}$. ∎
We shall eliminate the quadratic zeroth-order terms $\|f\|_{L^{2}(\Omega)}^{2}$ and $\|v\|_{L^{2}(\Omega)}^{2}$ left in Lemma 4.2 and Lemma 4.3 without treatment.
Lemma 4.4.
Let $\Omega=B_{R}\subset\mathbb{R}^{n}$ be a ball of radius $R>0$ with some $n\geq 5$. Then there exists a constant $C_{4}>0$ such that
[…]
$\quad+C_{4}(R^{2\kappa-n}+1)(B^{2}+m^{2}+M^{2}+1)\rho^{n-2\kappa}$
holds for all $(u,v)\in\mathcal{S}(m,M,B,\kappa)$ and $\rho\in(0,R)$, where $\mathcal{F}$ is given in (24).
Proof.
Lemma 4.1 and 4.2 entail there exists $C_{1}>0$ such that
-$\mathcal{F}$ $\leq 3\|\Delta v\|_{L^{2}(B_{\rho})}^{2}+3\|v\|_{L^{2}(B_{\rho})}^{2}+\|f\|_{L^{2}(B_{\rho})}^{2}+C_{1}(Bm+BM)\rho^{2-\kappa}+C_{1}$
holds for all $\rho\in(0,R)$. Furthermore, Lemma 4.3 warrants the existence of $C_{3}>0$ satisfying
$\begin{split}&-\frac{1}{24}\mathcal{F}\leq 12\rho^{2}\|\nabla f\|_{L^{2}(B_{\rho})}^{2}+C_{3}\|f\|_{L^{2}(B_{\rho})}^{2}+\sqrt{m}\rho\|g\|_{L^{2}(B_{\rho})}\\ &+\frac{9}{8}\|v\|_{L^{2}(B_{\rho})}^{2}-\frac{3}{4}\|\nabla v\|_{L^{2}(B_{\rho})}^{2}+C_{3}B^{2}\rho^{n-2\kappa}+C_{3}(Bm+BM)\rho^{2-\kappa}+2m+C_{1}\end{split}$ (43)
for all $\rho\in(0,R)$. Gagliardo-Nirenberg inequality [39] provides a constant $C_{gn}>0$ such that
$\|\psi\|_{L^{2}(\Omega)}^{2}\leq C_{gn}\|\nabla\psi\|_{L^{2}(\Omega)}^{2n/(n+2)}\|\psi\|_{L^{1}(\Omega)}^{4/(n+2)}+C_{gn}\|\psi\|_{L^{1}(\Omega)}^{2}$
holds for all $\psi\in W^{1,2}(\Omega)$. Recalling (25) we thus have
$\|f\|_{L^{2}(\Omega)}^{2}\leq C_{gn}B^{4/(n+2)}\|\nabla f\|_{L^{2}(\Omega)}^{2n/(n+2)}+C_{gn}B^{2},$ (44)
as well as
$\begin{split}\|v\|_{L^{2}(B_{\rho})}^{2}&\leq C_{gn}M^{4/(n+2)}\|\nabla v\|_{L^{2}(\Omega)}^{2n/(n+2)}+C_{gn}M^{2}\\ &\leq\frac{1}{2}\|\nabla v\|_{L^{2}(\Omega)}^{2}+2^{n}C_{gn}^{(n+2)/2}M^{2}+C_{gn}M^{2}\end{split}$ (45)
for all $\rho\in(0,R)$ by Young inequality. Invoking (26), we estimate
$\begin{split}\|\nabla v\|^{2}_{L^{2}(\Omega)}&=\|\nabla v\|_{L^{2}(B_{\rho})}^{2}+\|\nabla v\|^{2}_{L^{2}(\Omega\setminus B_{\rho})}\\ &\leq\|\nabla v\|_{L^{2}(B_{\rho})}^{2}+B^{2}\||x|^{1-\kappa}\|_{L^{2}(\Omega\setminus B_{\rho})}^{2}\\ &\leq\|\nabla v\|_{L^{2}(B_{\rho})}^{2}+\omega_{n}B^{2}R^{2}\rho^{n-2\kappa}\end{split}$ (46)
for all $\rho\in(0,R)$. Inserting (44)–(46) into (43) yields
$\begin{split}-\frac{1}{24}\mathcal{F}&\leq 12\rho^{2}\|\nabla f\|_{L^{2}(\Omega)}^{2}+C_{3}C_{gn}B^{4/(n+2)}\|\nabla f\|_{L^{2}(\Omega)}^{2n/(n+2)}+\sqrt{m}\rho\|g\|_{L^{2}(\Omega)}+C(\rho)\end{split}$
for all $\rho\in(0,R)$, where
$C(\rho)$ $:=C_{3}C_{gn}B^{2}+2^{n+1}C_{gn}^{(n+2)/2}M^{2}+2C_{gn}M^{2}+\omega_{n}B^{2}R^{2}\rho^{n-2\kappa}$
$\quad+C_{3}B^{2}\rho^{n-2\kappa}+C_{3}(Bm+BM)\rho^{2-\kappa}+2m+C_{1}$
$\>\leq(C_{3}C_{gn}+2^{n+1}C_{gn}^{(n+2)/2}+2C_{gn}+1)(B^{2}+M^{2}+m^{2}+1)$
$\quad+(\omega_{n}R^{2}+C_{3}+C_{3}R^{\kappa-n+2}+C_{1}R^{2\kappa-n})(B^{2}+M^{2}+m^{2}+1)\rho^{n-2\kappa}.$
Taking
$\begin{split}C_{4}&:=C_{3}C_{gn}+2^{n+1}C_{gn}^{(n+2)/2}+2C_{gn}+1+\omega_{n}+2C_{3}+C_{1},\end{split}$
we finish the proof. ∎
Here we shall choose $\rho$ appropriately, following the strategy of Winkler [6], to establish a lower bound of the energy $\mathcal{F}$ by a sublinear power of the dissipation rate $\mathcal{D}$.
Lemma 4.5.
Let $\Omega=B_{R}\subset\mathbb{R}^{n}$ be a ball of radius $R>0$ with some $n\geq 5$. Then there exist $\theta\in(0,1)$ and $C(m,M,B,\kappa)>0$ such that
$\mathcal{F}\leq C(m,M,B,\kappa)(\mathcal{D}^{\theta}+1)$ (47)
holds for all $(u,v)\in\mathcal{S}(m,M,B,\kappa)$, where $\mathcal{F}$ and $\mathcal{D}$ are specified as in (24).
Proof.
Put
$\rho:=\min\left\{\frac{R}{2},\|\nabla f\|_{L^{2}(\Omega)}^{-\frac{n}{(n+2)(2\kappa-n)}}\right\},$
and let
$\theta:=\begin{cases}\frac{n}{n+2}>1-\frac{n}{(n+2)(2\kappa-n)}>0,&\text{if }\kappa\in\big(n-2,\frac{3n}{4}\big),\\ 1-\frac{n}{(n+2)(2\kappa-n)}\geq\frac{n}{n+2},&\text{if }\kappa\in\big[\frac{3n}{4},\infty\big).\end{cases}$
We may infer the existence of $C(m,M,B,\kappa)>0$ from Lemma 4.4 such that
-$\mathcal{F}$ $\leq C(m,M,B,\kappa)(\rho^{2}\|\nabla f\|_{L^{2}(\Omega)}^{2}+\|\nabla f\|_{L^{2}(\Omega)}^{2n/(n+2)}+\|g\|_{L^{2}(\Omega)}+\rho^{n-2\kappa})$
holds for all $(u,v)\in\mathcal{S}(m,M,B,\kappa)$. If $\rho<R/2$, then
$\begin{split}-\mathcal{F}&\leq C(m,M,B,\kappa)(\rho^{2}\|\nabla f\|_{L^{2}(\Omega)}^{2}+\|\nabla f\|_{L^{2}(\Omega)}^{2n/(n+2)}+\|g\|_{L^{2}(\Omega)}+\|\nabla f\|_{L^{2}(\Omega)}^{n/(n+2)})\\ &\leq C(m,M,B,\kappa)\left(3\|\nabla f\|_{L^{2}(\Omega)}^{2\theta}+\|g\|_{L^{2}(\Omega)}^{2\theta}+4\right)\\ &\leq C(m,M,B,\kappa)(\mathcal{D}^{\theta}+1),\end{split}$ (48)
by Young inequality and monotonicity of the mapping $0,\infty)\ni\eta\mapsto\eta^{\theta}\in[0,\infty)$. On the other hand, if $\rho=R/2$, then we can estimate analogously
$\begin{split}-\mathcal{F}&\leq C(m,M,B,\kappa)\left(\rho^{2}\|\nabla f\|_{L^{2}(\Omega)}^{2}+\|\nabla f\|_{L^{2}(\Omega)}^{2n/(n+2)}+\|g\|_{L^{2}(\Omega)}+\frac{2^{2\kappa-n}}{R^{2\kappa-n}}\right)\\ &\leq C(m,M,B,\kappa)\left(2\|\nabla f\|_{L^{2}(\Omega)}^{2\theta}+\|g\|_{L^{2}(\Omega)}^{2\theta}+3+\frac{2^{2\kappa-n}}{R^{2\kappa-n}}\right)\\ &\leq C(m,M,B,\kappa)(\mathcal{D}^{\theta}+1).\end{split}$ (49)
Hence, (47) follows from (48) and (49). ∎
5 Finite-time blowup enforced by low initial energy
This section is devoted to the proof of Theorem 1.1. For initial data $(u_{0},v_{0})$ with low energy $\mathcal{F}(u_{0},v_{0})$, we may deduce a superlinear differential inequality for the mapping $t\mapsto-\mathcal{F}(u(\cdot,t),v(\cdot,t))$ from (47), which implies that $(u,v)$ cannot exist globally.
Lemma 5.1.
Assume $\Omega=B_{R}\subset\mathbb{R}^{n}$ for some $n\geq 5$ and $R>0$. Let $m>0$, $A>0$ and $\kappa>n-2$ be given. Then there exist positive constants $K(m,A,\kappa)$ and $T(m,A,\kappa)$ such that given any $(u_{0},v_{0})$ from the set
[…]
for the corresponding classical solution with maximal existence time $T_{\max}(u_{0},v_{0})$ given by Proposition 2.1, one has $T_{\max}(u_{0},v_{0})\leq T(m,A,\kappa)$.
Proof.
Let us fix $c_{1}>0$ such that
$\|\psi\|_{L^{1}(\Omega)}\leq c_{1}\|\psi\|_{W^{2,2}(\Omega)}\quad\text{for all }\psi\in W^{2,2}(\Omega).$
Lemma 2.3 and Proposition 3.1 give $c_{2}>0$, $c_{3}>0$ and $c_{4}=c_{4}(\kappa)>0$ such that whenever $(u_{0},v_{0})$ complies with (6), the corresponding solution $(u,v,w)$ satisfies $L^{1}$ estimate
$\|v_{t}\|_{L^{1}(\Omega)}\leq c_{2}(m+\|v_{0}\|_{W^{2,2}(\Omega)})\quad\text{for all }t\in(0,T_{\max}(u_{0},v_{0})),$ (51)
as well as pointwise estimates
$\||x|^{n-1}w\|_{W^{1,\infty}(\Omega)}\leq c_{3}m\quad\text{for all }t\in(0,T_{\max}(u_{0},v_{0})),$ (52)
and
$\||x|^{\kappa-1}v\|_{W^{1,\infty}(\Omega)}\leq c_{4}(m+\|v_{0}\|_{W^{2,2}(\Omega)})$ (53)
for all $t\in(0,T_{\max}(u_{0},v_{0}))$. Now writing
$B$ $:=\max\{c_{1}A,m,c_{3}m\}\quad\text{and}\quad M:=\max\{c_{2}(m+A),c_{4}(m+A)\},$
we invoke Lemma 4.5 to obtain $c_{5}=C(m,M,B,\kappa)>0$ and $\theta\in(0,1)$ such that
[…]
We will show the desired conclusion holds if we define
$K(m,A,\kappa):=c_{5}$ (55)
and
$T(m,A,\kappa):=\frac{\theta c_{5}^{\frac{1}{\theta}}}{(1-\theta)(-\mathcal{F}(u_{0},v_{0})-c_{5})^{\frac{1-\theta}{\theta}}}.$
Indeed, we know from Proposition 2.1 that for any given $(u_{0},v_{0})\in\widetilde{\mathcal{B}}(m,A,\kappa)$, the corresponding solution $(u,v,w)$ is smooth and radially symmetric with $u>0$ and $v>0$ in $\overline{\Omega}\times[0,T_{\max}(u_{0},v_{0}))$. We infer from (51)–(53) that
$\|v_{t}\|_{L^{1}(\Omega)}\leq c_{2}(m+A)\leq M\quad\text{for all }t\in(0,T_{\max}(u_{0},v_{0})),$
as well as pointwise estimates
$\||x|^{n-1}w\|_{W^{1,\infty}(\Omega)}\leq c_{3}m\leq B\quad\text{for all }t\in(0,T_{\max}(u_{0},v_{0})),$
and
$\||x|^{\kappa-1}v\|_{W^{1,\infty}(\Omega)}\leq c_{4}(m+A)\leq M$
for all $t\in(0,T_{\max}(u_{0},v_{0}))$. Since moreover $\int_{\Omega}u(x,t)=\int_{\Omega}u_{0}=m$ and
$\int_{\Omega}v(x,t)\leq\max\left\{\int_{\Omega}v_{0},\int_{\Omega}u_{0}\right\}\leq\max\{c_{1}A,m\}\leq B$
for all $t\in(0,T_{\max}(u_{0},v_{0}))$ by (12) and (13), it follows that $(u,v)\in\mathcal{S}(m,M,B,\kappa)$ for all $t\in(0,T_{\max}(u_{0},v_{0}))$ and thus (54) may be applied to $(\tilde{u},\tilde{v})=(u(\cdot,t),v(\cdot,t))$ for all $t\in(0,T_{\max}(u_{0},v_{0}))$. From (10) and (55), we see that
$\mathcal{F}(u(\cdot,t),v(\cdot,t))\leq\mathcal{F}(u_{0},v_{0})<-K(m,A,\kappa)=-c_{5}\quad\text{for all }t\in(0,T_{\max}(u_{0},v_{0})),$
and thus we may invert (54) to get
[…]
We shall see $T_{\max}(u_{0},v_{0})\leq T(m,A,\kappa)$. Otherwise a direct comparison argument by solving (56) gives
$(-\mathcal{F}(u,v)-c_{5})^{\frac{1-\theta}{\theta}}\geq\frac{\theta c_{5}^{\frac{1}{\theta}}}{(1-\theta)\left(T(m,A,\kappa)-t\right)}$
for all $t\in(0,T(m,A,\kappa))$. This warrants
$\mathcal{F}(u,v)\to-\infty\quad\text{as }t\nearrow T(m,A,\kappa),$
which is incompatible with Proposition 2.1. We thus verify $T_{\max}(u_{0},v_{0})\leq T(m,A,\kappa)$ as desired. ∎
Now we are in a position to show Theorem 1.1.
Proof of Theorem 1.1.
We only need to fix an arbitrary $\kappa>n-2$ and apply Lemma 5.1 to find that the conclusion holds if we let $K(m,A):=K(m,A,\kappa)$ and $T(m,A):=T(m,A,\kappa)$ with $K(m,A,\kappa)$ and $T(m,A,\kappa)$ given by Lemma 5.1. ∎
6 Initial data with large negative energy
This section is devoted to the existence of initial data with low energy.
Lemma 6.1.
Let $n\geq 5$ and $\Omega\subset\mathbb{R}^{n}$ be a bounded and smooth domain. Then for any choice of function pair $(u_{0},v_{0})$ satisfying (6), one can find a sequence
$\{(u_{k},v_{k})\in L^{\infty}(\Omega)\times W^{2,\infty}(\Omega)\}_{k\in\mathbb{N}}$
of function couples such that $(u_{k},v_{k})$ satisfies (6) for all $k\in\mathbb{N}$, that
$\int_{\Omega}u_{k}=\int_{\Omega}u_{0}\quad\text{for all }k\in\mathbb{N},$ (57)
that
[…]
that
[…]
and that
$\mathcal{F}(u_{k},v_{k})\to-\infty,\quad\text{as }k\to 0,$ (60)
where $\mathcal{F}$ is taken from (11).
Proof.
Without loss of generality, we may assume that $B_{R}\subset\Omega$ for some $R>0$. Let $\phi\in C_{0}^{\infty}(\mathbb{R}^{n})$ be a radially symmetric and nonnegative function with $\int_{\mathbb{R}^{n}}\phi=1$, which is compactly supported in $B_{1}$. Fixing $\gamma>0$, we define for $k\in\mathbb{N}$,
$u_{k}(x):=u_{0}(x)+\left(\ln\frac{1}{\eta}\right)^{2\gamma}\eta^{-\frac{n}{2}-2}\phi\left(\frac{x}{\eta}\right)-\left(\ln\frac{1}{\eta}\right)^{2\gamma}\eta^{\frac{n}{2}-2}\cdot\frac{u_{0}(x)}{\int_{\Omega}u_{0}},\quad x\in\overline{\Omega},$
and
$v_{k}(x):=v_{0}(x)+\left(\ln\frac{1}{\eta}\right)^{-\gamma}\eta^{2-\frac{n}{2}}\phi\left(\frac{x}{\eta}\right),\quad x\in\overline{\Omega},$
where
$\eta=\eta(k):=\frac{\eta_{\star}}{k+1},\quad k\in\mathbb{N},$
and $\eta_{\star}\in(0,1)\cap(0,R)$ is fixed such that
$\eta^{\frac{n}{2}-2}\left(\ln\frac{1}{\eta}\right)^{2\gamma}<\int_{\Omega}u_{0}\quad\text{for all }k\in\mathbb{N}.$ (61)
A direct computation can verify the mass identity (57). Since
$\|v_{k}-v_{0}\|_{W^{2,2}(\Omega)}\leq\left(\ln\frac{1}{\eta}\right)^{-\gamma}\|\phi\|_{W^{2,2}(B_{1})}$
are valid for all $k\in\mathbb{N}$, $\gamma>0$ ensures (58). We estimate by triangle inequality,
$\begin{split}\|u_{k}-u_{0}\|_{L^{p}(\Omega)}&\leq\left(\ln\frac{1}{\eta}\right)^{2\gamma}\eta^{-2-\frac{n}{2}+\frac{n}{p}}\|\phi\|_{L^{p}(B_{1})}\\ &\quad+\left(\ln\frac{1}{\eta}\right)^{2\gamma}\eta^{\frac{n}{2}-2}\cdot\frac{\|u_{0}\|_{L^{p}(\Omega)}}{\int_{\Omega}u_{0}}\quad\text{for all }k\in\mathbb{N}\text{ and }p\in[1,\infty).\end{split}$
Then (59) follows due to 2n/2+n/p>0-2-n/2+n/p>0 and $n>4$. To verify (60), we evaluate each term from $\mathcal{F}$. Young inequality entails
$\int_{\Omega}|v_{k}|^{2}\leq 2\|v_{0}\|_{L^{2}(\Omega)}^{2}+2\left(\ln\frac{1}{\eta}\right)^{-2\gamma}\eta^{4}\|\phi\|_{L^{2}(B_{1})}^{2}\quad\text{for all }k\in\mathbb{N}$ (62)
and
$\int_{\Omega}|\Delta v_{k}|^{2}\leq 2\|\Delta v_{0}\|_{L^{2}(\Omega)}^{2}+2\left(\ln\frac{1}{\eta}\right)^{-2\gamma}\|\Delta\phi\|_{L^{2}(B_{1})}^{2}\quad\text{for all }k\in\mathbb{N}.$ (63)
Thanks to (61), $u_{k}$ is nonnegative for all $k\in\mathbb{N}$, thus we can drop the nonnegative summands and estimate from below
$\begin{split}\int_{\Omega}u_{k}v_{k}&\geq\left(\ln\frac{1}{\eta}\right)^{\gamma}\int_{\Omega}\phi^{2}\left(\frac{x}{\eta}\right)\eta^{-n}\;\mathrm{d}x\\ &=\left(\ln\frac{1}{\eta}\right)^{\gamma}\|\phi\|_{L^{2}(B_{1})}^{2}\quad\text{for all }k\in\mathbb{N}.\end{split}$ (64)
Noting that
$\xi^{q}-\xi\geq(q-1)\xi\ln\xi\quad\text{for all }\xi\in(0,\infty)\text{ and }q\in(1,\infty),$
we compute
$\begin{split}\int_{\Omega}u_{k}\ln u_{k}&\leq\int_{\Omega}\frac{u_{k}^{q}-u_{k}}{q-1}\leq\int_{\Omega}\frac{u_{k}^{q}}{q-1}\\ &\leq\frac{2^{q-1}}{q-1}\bigg(\|u_{0}\|_{L^{q}(\Omega)}^{q}+\Big(\ln\frac{1}{\eta}\Big)^{2q\gamma}\eta^{-\frac{nq}{2}-2q+n}\|\phi\|_{L^{q}(B_{1})}^{q}\bigg)\end{split}$ (65)
for all $k\in\mathbb{N}$ and $q\in(1,\infty)$. Collecting (62)–(65) gives
$\begin{split}\mathcal{F}(u_{k},v_{k})&=\int_{\Omega}(u_{k}\log u_{k}-u_{k}v_{k})+\frac{1}{2}\int_{\Omega}|(-\Delta+1)v_{k}|^{2}\\ &\leq-\left(\ln\frac{1}{\eta}\right)^{\gamma}\|\phi\|_{L^{2}(B_{1})}^{2}+2\left(\ln\frac{1}{\eta}\right)^{-2\gamma}(\|\Delta\phi\|_{L^{2}(B_{1})}^{2}+\|\phi\|_{L^{2}(B_{1})}^{2})\\ &\quad+\frac{2^{q-1}}{q-1}\bigg(\ln\frac{1}{\eta}\bigg)^{2q\gamma}\eta^{-\frac{nq}{2}-2q+n}\|\phi\|_{L^{q}(B_{1})}^{q}\\ &\quad+2\|\Delta v_{0}\|_{L^{2}(\Omega)}^{2}+2\|v_{0}\|_{L^{2}(\Omega)}^{2}+\frac{2^{q-1}}{q-1}\|u_{0}\|_{L^{q}(\Omega)}^{q}\end{split}$
for all $k\in\mathbb{N}$ and
$q\in\bigg(1,\frac{2n}{n+4}\bigg).$
In view of $\gamma>0$, we have $\mathcal{F}(u_{k},v_{k})\to$-$\infty$ as $k\to\infty$ and complete the proof. ∎
Now we are in a position to show Theorem 1.2.
Proof of Theorem 1.2.
Let initial datum $(u_{0},v_{0})$ be specified as in Theorem 1.2. Then the sequence $\{(u_{k},v_{k})\}_{k\in\mathbb{N}}$ of positive function pairs constructed in Lemma 6.1 clearly satisfies the desired properties according to (57)–(60). ∎
[…]
*Acknowledgements The authors thank the anonymous referees for their helpful comments and suggestions, which greatly improve the presentation of our paper.
[…]
*Funding The first author has been supported by “the Fundamental Research Funds for the Central Universities” (No. B250201215). The second author has been supported in part by National Natural Science Foundation of China (No. 12271092, No. 11671079) and the Jiangsu Provincial Scientific Research Center of Applied Mathematics under Grant No. BK20233002.
[…]
*Conflict of interest
The authors have not disclosed any conflict of interest.